### v2

# Conditional existence of maximizers for the Tomas-Stein inequality for the sphere

Abstract.
The Tomas-Stein inequality for a compact subset $\Gamma$ of the sphere $S^{d}$ states that the mapping $f\mapsto\widehat{f\sigma}$ is bounded from $L^{2}(\Gamma,\sigma)$ to $L^{2+4/d}({\mathbb{R}}^{d+1})$. Then conditional on a strict comparison between the best constants for the sphere and for the Strichartz inequality for the Schrödinger equations, we prove that there exist functions which extremize this inequality, and any extremising sequence has a subsequence which converges to an extremizer. The method is based on the refined Tomas-Stein inequality for the sphere and the profile decompositions. The key ingredient to establish orthogonality in profile decompositions is that we use Tao’s sharp bilinear restriction theorem for the paraboloids beyond the Tomas-Stein range. Similar results have been previously established by Frank, Lieb and Sabin [14], where they used the method of the missing mass.
1. Introduction
The Tomas-Stein’s inequality for the sphere or the adjoint Fourier restriction inequality for a compact subset of the sphere $\Gamma\subset S^{d}$ asserts that
(1) $\|\widehat{f\sigma}\|_{L^{2+4/d}({\mathbb{R}}^{d+1})}\leq\mathcal{R}\|f\|_{L^{2}(\Gamma,\sigma)}.$
See e.g. [31]. Here the constant $\mathcal{R}>0$ is defined to be the optimal constant
(2) $\begin{split}\mathcal{R}:=\sup\{\|\widehat{f\sigma}\|_{L^{2+4/d}({\mathbb{R}}^{d+1})}:\|f\|_{L^{2}(\Gamma,\sigma)}=1\},\\ \Gamma:=\{x=(x^{\prime},x_{d+1}):\,x\in S^{d},\,|x^{\prime}|\leq\frac{1}{2}\}.\end{split}$
and $\sigma$ denotes the surface measure on $\Gamma\subset S^{d}$, and the Fourier transform is defined by
(3) $\widehat{f\sigma}(\xi):=\frac{1}{(2\pi)^{\frac{d+1}{2}}}\int_{\Gamma}e^{-i\xi\cdot x}f(x)d\sigma(x).$
Definition 1.1.
A function $f\in L^{2}(\Gamma,\sigma)$ is said to be an extremal for (1) if $f\neq 0$ a. e., and
(4) $\|\widehat{f\sigma}\|_{L^{2+4/d}({\mathbb{R}}^{d+1})}=\mathcal{R}\|f\|_{L^{2}(\Gamma,\sigma)}.$
An extremising sequence for the inequality (1) is a sequence $\{f_{\nu}\}\in L^{2}(\Gamma,\sigma)$ satisfies $\|f_{\nu}\|_{L^{2}(\Gamma,\sigma)}=1$ and $\lim_{\nu\to\infty}\|\widehat{f_{\nu}\sigma}\|_{L^{2+4/d}({\mathbb{R}}^{d+1})}=\mathcal{R}$. An extremising sequence is said to be precompact if any subsequence has a sub-subsequence which is Cauchy in $L^{2}(\Gamma,\sigma)$.
To prove the existence of extermizers for the Tomas-Stein inequality, we need to establish the compactness of the complex $L^{2}(S^{d})$ extremizing sequence. The noncompactness of the Tomas-Stein inequality is due to the modulation symmetry: $x\in S^{d}\mapsto e^{x\cdot\xi}$ given $\xi\in\mathbb{C}^{d+1}$, and the approximate scaling symmetry. The symmetry $f\mapsto f\cdot e^{ix\cdot\xi}$ merits further discussion and is closely related to the approximate symmetry. For the paraboloid $\mathbb{P}^{d}:=\{(y_{1},\cdots,y_{d},y_{d+1}):y_{d+1}=\frac{1}{2}(y_{1}^{2}+\cdots y_{d+1}^{2})\}$, the analogy of the unimodular exponential $e^{ix\cdot\xi}$are quadratic exponentials $e^{i\eta\cdot x+it\tau|x|^{2}}$ with $(\eta,\tau)\in\mathbb{R}^{d+1}$, where $\xi,\eta\in\mathbb{R}^{d}$ ranges over a $d$-dimensional space. To see the analogy, consider a small neighborhood of $(0,\cdots,0,1)\in S^{d}$, equipped with coordinates $x^{\prime}\in\mathbb{R}^{d}$ so that $x=(x^{\prime},x_{d+1})$. Then for $\xi=(0,\cdots,0,\lambda)$, $e^{ix\cdot\xi}=\exp\bigl(i\lambda(1-\frac{1}{2}|x^{\prime}|^{2}+O(|x^{\prime}|^{4}))\bigr)$ for small $x^{\prime}$; thus for small $x^{\prime}$, one has essentially quadratic oscillation. The presence of these symmetries among the extremizers for $\mathbb{P}^{d}$ implies that, in the language of the concentration-compactness theory [17, 11, 15, 27], an extremizer $f$ can be tight at a scale $r$, and $\hat{f}$ can simultaneously be tight at a scale $\hat{r}$, with the product $r\cdot\hat{r}$ arbitrarily large and thus loss of compactness is manifested, see Equation (67).
We define the optimal constant for the Strichartz inequality for the Schrödinger equations.
(5) $\mathcal{R}_{\mathbf{P}}=(2\pi)^{-\frac{d+2}{d}}\dfrac{\int_{\mathbb{R}^{d+1}}|e^{it\Delta/2}f(x)|^{2+\frac{4}{d}}dxdt}{\|f\|^{2+\frac{d}{4}}_{L^{2}(\mathbb{R}^{d})}},$
Theorem 1.2.
Assume that $\mathcal{R}>\mathcal{R}_{\mathbf{P}}$. Then there exists an extremal function $f\in L^{2}(\Gamma,\sigma)$ for (1) by showing that any extremising sequence $\{f_{\nu}\}$ is precompact in $L^{2}(\Gamma,\sigma)$. This extremizer can be chosen to be a smooth function.
Remark 1.3.
In [14], Frank, Lieb and Sabin assume that Gaussian functions are the extremizers optimizing (5) and prove that $\mathcal{R}>\mathcal{R}_{\mathbf{P}}$ by a perturbative analysis. Similar analysis has previously appeared in [9, 10].
Let $\mathbb{P}^{d}$ be the paraboloid introduced above. Let $\sigma_{\mathbb{P}^{d}}$ be the measure $d\sigma_{\mathbb{P}^{d}}=dx_{1}\cdots dx_{d}$ on $\mathbb{P}^{d}$. The mapping $f\mapsto\widehat{f\sigma_{\mathbb{P}^{d}}}$ is likewise bounded from $L^{2}(\mathbb{P}^{d},\sigma_{\mathbb{P}^{d}})$ to $L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})$. Using $\mathcal{R}_{\mathbf{P}}$,
$e^{it\Delta/2}f(t,x)$ $:=\int_{\mathbb{R}^{d}}e^{ix\cdot y-i\frac{t|y|^{2}}{2}}f(y)dy.$
$\|e^{it\Delta/2}f\|_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}$ $\leq(2\pi)^{1/2}\mathcal{R}_{\mathbf{P}}\|f\|_{L^{2}(\mathbb{R}^{d})}.$
When $d=1,2$, Foschi [11] has proved that extremals exist for this inequality, and moreover that every radial Gaussian $f(x^{\prime},x_{3})=e^{-c|x^{\prime}|^{2}}$ is an extremal, where $x^{\prime}=(x_{1},x_{2})\in\mathbb{R}^{2}$. Alternative proofs are given by Hundertmark and Zharnitisky [15] and Bennett, Bez, Carbery and Hundertmark [3]. The simple relation $\mathcal{R}\geq\mathcal{R}_{\mathbf{P}}$ is of significance. The relation follows from examination of a suitable sequence of trial functions $f_{\nu}$, such that $f_{\nu}(x)^{2}$ converges weakly to a Dirac mass on $S^{d}$, and $f_{\nu}$ is approximately a Gaussian in suitably rescaled coordinates depending on $\nu$. It is essential for this comparison that $\mathbb{P}^{d}$ has the same curvature at 0 as $S^{d}$, which explains the factors of $\frac{1}{2}$ in the definition of $\mathbb{P}^{d}$.
Kunze [17] is the first to discuss the existence of extremizers for the Strichartz/Fourier restriction inequalities and proved the existence of extremizers for the parabola in $\mathbb{R}^{2}$, and showed that any nonnegative extremizing sequence is precompact modulo the action of the natural symmetry group of the inequality. Several papers have subsequently dealt with related problems, in some cases determining all the extremizers explicitly [11, 15, 3], in other cases merely proving existence [27]. A powerful result which leads easily [27] to existence of extremizers is the profile decomposition [1, 2]. Of these works, the one closely related to ours is that of [1, 2, 27, 26]. One difficulty which we face is the lack of exact scaling symmetries. In some facets of the analysis this is merely a technical obstacle, but it is bound up with the most essential obstacle, which is the possibility that the optimal constant might be achieved only in a limit where $|f_{\nu}|^{2}$ tends to a Dirac mass on $\Gamma\subset S^{d}$.
Our analysis follows the general profile decomposition framework in [9, 1, 2, 27, 26]. One earlier example is that the author proved the existence of extermizers for the one-dimensional Tomas-Stein inequality [28]; which has precursors in [9] and [10]. In the latter two works, Christ and the first author proved existence of extremals for the Tomas-Stein inequality for $S^{2}$, and obtained some exact characterizations of nonnegative extremals, complex extremals and complex extremising sequences. The general idea is the profile decomposition or the concentration-compactness approach [19, 20, 21, 22]. Roughly speaking, the profile decomposition starts with Bourgain’s $X_{p}$-refinement of the Tomas-Stein inequality for the sphere and obtains a near-extremizer decomposition. Each piece will be a $L^{2}$-normalized cap on the sphere. The second decomposition of this cap to upgrade the weak convergence to strong convergence is to display the approximate scaling symmetry, which resembles the Schrödinger behavior. The key to prove the orthogonality of profiles is to use the Tao bilinear restriction estimates for elliptic surfaces, see Proposition 5.1. The novelty of our work is that we apply this estimate in the concentration-compactness theory of the maximizer problem for the Fourier restriction inequalities to establishing orthogonality of profiles. The analysis of the extremizer problem for the Tomas-Stein inequality for the two dimensional sphere in [9] has extended to other manifolds [25, 8], and other norms (nonendpoint or mixed-norms) in [6, 7]. See the references therein. The current work aims to address the same question in [9] in higher dimensions. Note again that Frank, Lieb and Sabin answered this question in [14], though our method of profile decompositions is different from theirs.
In [9], Christ and the first author have proved that constant functions are local extremizers to the Tomas-Stein inequality for $S^{2}$. The key ingredients are the use of spherical harmonics [30, 36]. This idea was later used by Foschi [13] to establish that constants are global maximizers for this inequality. In this preprint, we are not able to find the exact maximizers, though we believe that they are constant functions from the information of the existence proof in Section 7. The next natural step would be establishing that constants are local extremizers.
The paper is divided into three parts: the first part focuses on establishing the refined Tomas-Stein estimates in the spirit of Bourgain in Section 2, Lemma 2.2. In Sections 3, 4, 5, 6, we establish the theorem of profile decompositions for a sequence of $L^{2}(\Gamma,\sigma)$-extremising sequence, Propositions 6.1 and 6.2. In Section 7, we establish the conditional existence of extremizers for the Tomas-Stein inequality for $\Gamma\subset S^{d}$ in dimensions $d\geq 2$, Theorem 1.2.
Acknowledgements
The first author would like to thank Michael Christ for many helpful discussions. This work grows out of an early discussion with him in 2010. The authors have used the Chatgpt and overleaf to edit the figures in the paper.
2. The refined Tomas-Stein estimate
Definition 2.1.
The cap $\mathcal{C}=\mathcal{C}(z,r)$ with center $z\in S^{d}$ and radius $r\in(0,1]$ is the set of all points $y\in S^{d}$ which lie in the same hemisphere as $z$ and are centered at $z$, and which satisfy $\pi_{H_{z}}(y)<r$, where the subspace $H_{z}\subset{\mathbb{R}}^{d+1}$ is the orthogonal complement of $z$ and $\pi_{H_{z}}$ denotes the orthogonal projection onto $H_{z}$.
The following lemma on the refinement of the Tomas-Stein inequality is taken from [2, Theroem 1.3].
Lemma 2.2.
For $f\in L^{2}(\Gamma,\sigma)$. There exists $\alpha\in(0,1)$ such that
(6) $\|\widehat{f\sigma}\|_{2+4/d}\leq C\left(\sup_{\mathcal{C}}\frac{1}{|\mathcal{C}|^{1/2}}\int_{\mathcal{C}}|f|d\sigma\right)^{\alpha}\|f\|^{1-\alpha}_{L^{2}(\Gamma,\sigma)},$
where $\mathcal{C}$ denotes a cap on $S^{d}$.
To prove this lemma, we will use the Tao’s blinear restriction estimate for paraboloids [33]. Recently Frank, Lieb and Sabin directly prove this estimate in [14, Proposition 5.1] using similar ideas. We will provide a proof for completeness, though the outline is essentially contained in Begout-Vargas [2]. Namely we first prove a Bourgain’s $X_{p}$ estimate, Theorem 2.3. see the precursors in [4, 23], and then deduce this estimate from it. We remark this $X_{p,q}$ estimate of the Strichartz norm of the sphere restriction operator $\widehat{f\sigma}$ is the first time of proof in the literature.
For each integer $k\geq 0$ choose a maximal subset $\{z_{k}^{j}\}\subset S^{d}$ satisfying $|z_{k}^{j}-z_{k}^{i}|>2^{-k}$ for all $i\neq j$. Then for any $x\in S^{d}$ there exists $z_{k}^{i}$ such that $|x-z_{k}^{i}|\leq 2^{-k}$; otherwise $x$ could be adjoined to $z_{k}^{i}$ for this $z_{k}^{i}$, contradicting maximality. Therefore the caps $\mathcal{C}_{k}^{j}=(\mathcal{C}_{k}^{j},2^{-k+1})$ cover $S^{d}$ for each $k$, and there exists $C<\infty$ such that for any $k$, no point of $S^{d}$ belongs to more than $C$ of the caps $\mathcal{C}_{k}^{j}$. The constant $C$ is independent of $k$.
For $p\in[1,\infty)$, the $X_{p}$ norm is defined by
(7) $\|f\|_{X_{p,q}}=\left(\sum_{k}\sum_{j}|\mathcal{C}_{k}^{j}|^{q/2}\big(\frac{1}{|\mathcal{C}_{k}^{j}|}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\big)^{q/p}\right)^{1/q}.$
Moyua, Vargas, and Vega [24] have proved
Theorem A.
There exist $0<C<\infty$ and $p\in(1,2)$ such that for any $f\in L^{2}(S^{2})$,
(8) $\|{\widehat{f\sigma}}\|_{L^{4}(\mathbb{R}^{3})}\leq C\|f\|_{X_{p}}.$
We generalize this lemma to arbitrary dimensions.
Theorem 2.3.
Let $q=\frac{2(d+2)}{d}$ and $p<2$ be such that $\frac{1}{p^{\prime}}>\frac{d+3}{d+1}\frac{1}{q}$. For every function $f\in X_{p,q}$, we have
$\|\widehat{f\sigma}\|_{L^{q}(\mathbb{R}^{d+1})}\leq\|f\|_{X_{p,q}}.$
We follow the outline in [2] and give the details of the proof for the sake of completeness. Our main tool is the following sharp bilinear restriction estimate proved by Tao [33, Remark 3]. The precursor of this estimate and the sharp bilinear restriction estimate for the cone is proved by Wolff [35].
We write the upper quarter sphere as $\Gamma\subset S^{d},\,\Gamma=\{(x,\sqrt{1-|x|^{2}}):\,|x|\leq\frac{1}{2}\}$. Let $\phi(x)=1-\sqrt{1-|x|^{2}}$. For $f\in L^{2}(\Gamma,\sigma)$,
$e^{-it}\widehat{f\sigma}(t,x)=\int e^{ix\cdot\xi+it\phi(\xi)}f(\xi,\sqrt{1-|\xi|^{2}})\frac{d\xi}{\sqrt{1-|\xi|^{2}}}=\int e^{ix\cdot\xi+it\phi(\xi)}\frac{f(\xi,\sqrt{1-|\xi|^{2}})}{\sqrt{1-|\xi|^{2}}}d\xi.$
Since we are on the upper quarter sphere, the Jacobian factor $\sqrt{1-|\xi|^{2}}\sim 1$; we regard $\frac{f(\xi,\sqrt{1-|\xi|^{2}})}{\sqrt{1-|\xi|^{2}}}$ as a $d$-dimensional function on $\mathbb{R}^{d}$. We still denote it by $f$ in this section.
Theorem B.
[33] Let $Q,Q^{\prime}$ be cubes of sidelength $1$ in $\mathbb{R}^{d}$ such that
$\inf\{d(x,y):\,x\in Q,y\in Q^{\prime}\}\sim 1$
and $f,g$ functions supported in $Q$ and $Q^{\prime}$ respectively. Then for all $r>\frac{d+3}{d+1}$ and $p\geq 2$, we have
$\|\widehat{f\sigma}\widehat{g\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C\|f\|_{L^{p}(\mathbb{R}^{d})}\|g\|_{L^{p}(\mathbb{R}^{d})},$
with a constant $C$ independent of $f,g,Q,Q^{\prime}$.
By interpolation with the trivial estimate
$\|\widehat{f\sigma}\widehat{g\sigma}\|_{L^{\infty}(\mathbb{R}^{d+1})}\leq C\|f\|_{L^{1}}\|g\|_{L^{1}}\leq C\|f\|_{L^{p}}\|g\|_{L^{p}}$
for all $p\geq 1$, we can obtain
Theorem C.
[33] Let $Q,Q^{\prime}$ be cubes of sidelength $1$ in $\mathbb{R}^{d}$ such that
$\inf\{d(x,y):\,x\in Q,y\in Q^{\prime}\}\sim 1$
and $f,g$ functions supported in $Q$ and $Q^{\prime}$ respectively. Then for all $r>\frac{d+3}{d+1}$ and $p$ such that $\frac{2}{p^{\prime}}>\frac{d+3}{d+1}\frac{1}{r}$, we have
$\|\widehat{f\sigma}\widehat{g\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C\|f\|_{L^{p}(\mathbb{R}^{d})}\|g\|_{L^{p}(\mathbb{R}^{d})},$
with a constant $C$ independent of $f,g,Q,Q^{\prime}$.
By a simple argument using Theorem C, Frank, Lieb and Sabin obtained the following theorem [14, Theorem A. 1].
Theorem D.
[14] Let $\tau,\tau^{\prime}$ be cubes of sidelength $2^{-j}$ in $\mathbb{R}^{d}$ such that
$\inf\{d(x,y):\,x\in\tau,y\in\tau^{\prime}\}\sim 2^{-j}$
and $f,g$ functions supported in $\tau$ and $\tau^{\prime}$ respectively. Then for all $r=\frac{d+2}{d}$ and $p$ such that $\frac{2}{p^{\prime}}>\frac{d+3}{d+1}\frac{1}{r}$, we have
$\|\widehat{f\sigma}\widehat{g\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C2^{jd\frac{2-p}{p}}\|f\|_{L^{p}(\mathbb{R}^{d})}\|g\|_{L^{p}(\mathbb{R}^{d})},$
with a constant $C$ independent of $f,g,\tau$ and $\tau^{\prime}$.
We will need to use the orthogonality of functions with disjoint supports. More precisely, the following Bernstein type lemma, a proof of which can be found, for instance, in Tao, Vargas and Vega [34, Lemma 6.1].
Theorem E.
Let $\{R_{k}\}_{k\in\mathbb{Z}}$ be a collection of rectangles in frequency space and $c>0$, such that the dilates $(1+c)R_{k}$ are almost disjoint $i.e.,\sum_{k}\chi_{(1+c)R_{k}}\leq C$, and suppose that $(f_{k})_{k\in\mathbb{Z}}$ is a collection of functions whose Fourier transforms are supported in $R_{k}$. Then for all $1\leq p\leq\infty$, we have
$\|\sum_{k\in\mathbb{Z}}f_{k}\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,c)\left(\|f_{k}\|^{p^{*}}_{L^{p}(\mathbb{R}^{d})}\right)^{\frac{1}{p^{*}}},$
where $p^{*}=\min(p,p^{\prime})$.
Now we begin the proof of Theorem 2.3 and then deduce Theorem 2.2 from it.
Proof.
We set $r=\frac{q}{2}=\frac{d+2}{d}$. We can assume that the support of $f$ is contained in the square $\{x\in\mathbb{R}^{d}:\,|x_{i}|\leq\frac{1}{2},1\leq i\leq d\}$. For each $j\in\mathbb{Z}$, we decompose $\mathbb{R}^{d}$ into dyadic cubes $\tau_{j}^{k}$ of sidelength $2^{-j}$. Given a dyadic cube $\tau_{j}^{k}$ we will say that is the “parent” of the $2^{d}$ dyadic cubes of sidelength $2^{-j-1}$ contained in it. We define that $\tau_{k}^{j}$ is “orthogonal” to $\tau_{k^{\prime}}^{j}$, denoted by $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$, if $\tau^{j}_{k}$ and $\tau^{j}_{k^{\prime}}$ are not adjacent but have adjacent parents. For each $j\geq 0$, we write $f=\sum f_{k}^{j}$ where $f_{k}^{j}=f\chi_{\tau_{k}^{j}}$. Denote by $\beta$ the diagonal of $\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\beta=\{(x,x):\,x\in\mathbb{R}^{d}\}$. We have the following Whitney type decomposition, as indicated in Figure 1, of $\mathbb{R}^{d}\times\mathbb{R}^{d}\setminus\beta$,
$1/2,1/2]^{d}\times[-1/2,1/2]^{d}\setminus\beta=\cup_{j}\cup_{k,k^{\prime}:\,\tau^{j}_{k}\sim\tau^{j}_{k^{\prime}}}\tau_{k}^{j}\times\tau^{j}_{k^{\prime}}.$
Figure 1. The Whitney type decomposition.
This has previously appeared in [34] and [16, Proposition 4.24]. Indeed, given $\xi\neq\xi^{\prime}$, there is a unique maximal pair $\xi\in\tau$ and $\xi^{\prime}\in\tau^{\prime}$ such that
$\tau\sim\tau^{\prime},\text{ and }|\tau|=|\tau^{\prime}|,\text{ and }\operatorname{dist}(\tau,\tau^{\prime})\geq\frac{1}{\sqrt{d}}\operatorname{diam}(\tau).$
Let $\mathcal{D}$ denote the collection of such pairs $\tau\sim\tau^{\prime}$ as $\xi\neq\xi^{\prime}$ varies over $\frac{1}{2},\frac{1}{2}]^{d}.$ According to this definition, we have
$\sum_{(\tau,\tau^{\prime})\in\mathcal{D}}\chi_{\tau}(\xi)\chi_{\tau^{\prime}}(\xi^{\prime})=1,\text{ for a.e. }(\xi,\xi^{\prime})\in[-\frac{1}{2},\frac{1}{2}]^{d}.$
Note that since $\tau$ and $\tau^{\prime}$ are maximal, $\operatorname{dist}(\tau,\tau^{\prime})\leq 4\operatorname{diam}(\tau).$ In addition, this shows that given $\tau$, there is a bounded number of $\tau^{\prime}$ such that $(\tau,\tau^{\prime})\in\mathcal{D}$. That is to say,
$\forall\tau,\#\{\tau^{\prime},\,\tau^{\prime}\sim\tau\}\lesssim 1.$
Note that this is different from the usual Whitney decomposition, see e.g. [29, Chapter 1, Theorem 3].
So we can decompose $\widehat{f\sigma}$ correspondingly.
$e^{-2it}\widehat{f\sigma}(t,x)\widehat{f\sigma}(t,x)$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}e^{i(x\cdot\xi+t\phi(\xi))}e^{i(x\cdot\eta+t\phi(\eta))}f(\xi)g(\eta)d\xi d\eta$
$=\sum_{j}\sum_{k}\sum_{k^{\prime}:\,\tau^{j}_{k}\sim\tau^{j}_{k^{\prime}}}\widehat{f_{k}^{j}\sigma}(t,x)\widehat{f_{k^{\prime}}^{j}\sigma}(t,x)$
[…]
$\operatorname{supp}\mathcal{F}_{t,x}\left(\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\right)\subset\tilde{\tau_{k}^{j}}+\tilde{\tau_{k^{\prime}}^{j}}=\{\bigl(\phi(\xi)+\phi(\xi^{\prime}):\,\xi+\xi^{\prime}\bigr):\,\xi\in\tau_{k}^{j},\xi^{\prime}\in\tau_{k^{\prime}}^{j}\}.$
Here $\mathcal{F}_{t,x}(f)(\tau,\xi)=(2\pi)^{-\frac{d+1}{2}}\int_{\mathbb{R}^{d+1}}e^{i\xi\cdot x+it\tau}f(\tau,\xi)d\tau d\xi$. The $\phi$ is the elliptic phase defined in Tao, Vargas, and Vega [34, Page 968] and Tao [33, Page 1381]. We follow the approach outlined in [14, Lemma A.2] or [16, Proposition 4.24]. The Taylor expansion leads to the formula
$\phi(\xi)+\phi(\xi^{\prime})$ $=2\phi(\frac{\xi+\xi^{\prime}}{2})+a(\xi,\xi^{\prime})|\xi-\xi^{\prime}|^{2},$
$\phi(\frac{\xi+\xi^{\prime}}{2})$ $=\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})+\frac{1}{2}\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\xi+\xi^{\prime}-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))$
$+b(\eta+\eta^{\prime},c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))|\xi+\xi^{\prime}-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})|^{2},$
where $c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ denotes the center of the cube $\tau_{k}^{j}+\tau_{k^{\prime}}^{j}$, for two functions $a$ and $b$ satisfying
$\frac{3}{8}\leq a(\xi,\xi^{\prime})\leq-\frac{1}{8},\,-\frac{3}{16}\leq b(\xi+\xi^{\prime},c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))\leq-\frac{1}{16},$
assuming that $\epsilon$ in defining the elliptic phase in Tao, Vargas, and Vega [34, Page 968] and Tao [33, Page 1381] is sufficiently small. Further the space-time Fourier transform
$\operatorname{supp}\mathcal{F}_{t,x}\left(\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\right)\subset R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}),$
where
$R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}):=\left\{(b,a):\,a\in\tau_{k}^{j}+\tau_{k^{\prime}}^{j},\,-2\leq\frac{b-2\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(a-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))}{\operatorname{diam}(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})^{2}}\leq-\frac{1}{32}\right\}.$
Note that in computing these bounds forming the set, we have used
$0\leq|a-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})|\leq\operatorname{diam}(\tau_{k}^{j}),\,\operatorname{diam}(\tau_{k}^{j})\leq|\xi-\xi^{\prime}|\leq 4\operatorname{diam}(\tau_{k}^{j}),\operatorname{diam}(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})=2\operatorname{diam}(\tau_{k}^{j}).$
Again by the Taylor expansion, we have
$\frac{3}{16}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq\phi(\frac{a}{2})-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot\frac{1}{2}(a-c(\tau_{k}^{j}+\tau_{k}^{j}))\leq-\frac{1}{16}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}.$
Therefore for $(b,a)\in R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$, noting that $0\leq|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|\leq\operatorname{diam}(\tau_{k}^{j})$,
$\begin{split}b-2\phi(\frac{a}{2})&=-2\left(\phi(\frac{a}{2})-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\right)\\ &+b+2\left(-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\right).\end{split}$
Then
$8\operatorname{diam}(\tau_{j}^{k})^{2}+\frac{1}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq b-2\phi(\frac{a}{2})\leq-\frac{1}{8}\operatorname{diam}(\tau_{k}^{j})^{2}+\frac{3}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}.$
That is to say,
$8\operatorname{diam}(\tau_{j}^{k})^{2}\leq b-2\phi(\frac{a}{2})-\frac{1}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2},\text{ and }b-2\phi(\frac{a}{2})-\frac{3}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq-\frac{1}{8}\operatorname{diam}(\tau_{k}^{j})^{2}.$
Therefore if two pair of close cubes $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$ and $\tau_{m}^{n}\sim\tau_{m^{\prime}}^{n}$ are such that $R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ and $R(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})$ intersect, they must have a similar diameter. The same holds for the dilates $(1+\alpha)R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ for some small $\alpha$ by the same argument. Indeed, we can take $\alpha<0.01$. Lastly if the diameters are in a finite number, the cubes are also in a finite number since their centers satisfy
$|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-c(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})|\leq|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|+|c(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})-a|\leq 2\operatorname{diam}(\tau_{j}^{k}).$
Then we see that
(9) $\sup_{b,a}\sum_{(\tau_{k}^{j},\tau_{k^{\prime}}^{j})\in\mathcal{D}}\chi_{(1+\alpha)R(c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}(b,a)\leq C_{d},$
for some constant $C_{d}>0$. Then
$\|\widehat{f\sigma}\|^{2}_{L^{2r}(\mathbb{R}^{d+1})}=\|\widehat{f\sigma}\widehat{f\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}=\|\sum_{\mathcal{D}}\widehat{f_{k}^{j}\sigma}\widehat{g_{k^{\prime}}^{j}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}.$
By Theorem E, since $r<2$, we have
$\|\sum_{\mathcal{D}}\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C_{d}\left(\sum_{\mathcal{D}}\|\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}.$
Now we use Theorem D to obtain
$\left(\sum_{\mathcal{D}}\|\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}$
$\leq C_{d,p}\left(\sum_{j}\sum_{k}\sum_{k^{\prime}}2^{jdr\frac{2-p}{p}}\|f_{j}^{k}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{k^{\prime}}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r},$
where $p,q,r$ are assumed in Theorem D. Now for each $(j,k)$ there are at most $2d$ indices $k^{\prime}$ such that $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$. Hence
$\left(\sum_{j}\sum_{k}\sum_{k^{\prime}}2^{ldr\frac{2-p}{p}}\|f_{k}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{k^{\prime}}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
$\leq C_{d}\left(\sum_{j}\sum_{l}2^{jdr\frac{2-p}{p}}\|f_{k}^{j}\|^{2r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
The last sum is recognized to be $\|f\|_{X_{p,q}}$ with $q=2r=\frac{2(d+2)}{d}$ and $\frac{1}{p^{\prime}}>\frac{d+3}{d+1}\frac{1}{q}$. Thus
$\|\widehat{f\sigma}\|_{L^{q}(\mathbb{R}^{d+1})}\leq\|f\|_{X_{p,q}}.$
Now we finish the proof of Theorem 2.3. ∎
By using this Bourgain’s $X_{p,q}$ estimate, we will prove Lemma 2.2.
Lemma 2.4.
For $p\in(1,2)$. Then there exists $\mu\in(0,\frac{1}{p})$ such that for any $f\in L^{2}(\Gamma,\sigma)$, we have
(10) $\|\widehat{f\sigma}\|_{L^{2+\frac{4}{d}}}\leq C\left(\sup_{(j,k)\in\mathbb{Z}\times\mathbb{Z}^{d}}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\mu}\|f\|^{1-\mu p}_{L^{2}(S^{d})}.$
where $C=C(p,q)$ and $\mu=\mu(p,q)$.
If this lemma were established, then Lemma 2.2 follows by a simple application of Hölder inequality. For its proof, we will follow [2, Theorem 1.3].
Proof.
Step 1. By homogeneity, we can assume that $\|f\|_{L^{2}(\Gamma,\sigma)}=1$. Then it suffices to show that for any function $f\in L^{2}(S^{d})$ such that $\|f\|_{L^{2}(\Gamma,\sigma)}=1$,
$\sum_{k}\sum_{j}|\mathcal{C}_{k}^{j}|^{q/2}\big(\frac{1}{|\mathcal{C}_{k}^{j}|}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\big)^{q/p}\leq C(p,q)\left(\sup_{j,k}\left(|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}}\bigl(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\bigr)^{1/p}\right)\right)^{\alpha},$
where $\alpha=\mu pq$ and where $\mu$ has to be determined. We take $\alpha$ and $\beta$ such that $\frac{2}{q}<\beta<1$, $\beta>\frac{p}{2}$ and $\alpha+q\beta=q$. Such $\alpha,\beta$ are possible if $\mu$ is chosen such that
$0<\mu<\min\{\frac{1}{p}-\frac{1}{2},\,\frac{2}{(d+2)p}\}.$
Then taking out the $L^{\infty}$ norm directly,
$\begin{split}&\sum_{k}\sum_{j}|\mathcal{C}_{k}^{j}|^{q/2}\big(\frac{1}{|\mathcal{C}_{k}^{j}|}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\big)^{q/p}=\sum_{k}\sum_{j}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}q}\big(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\big)^{q/p}\\ &\leq C(p,q)\left(\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\frac{\beta q}{p}}\right)\left(\sup_{j,k}\left(|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}}\bigl(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\bigr)^{1/p}\right)\right)^{\alpha}\end{split}$
We set $\mu=\frac{\alpha}{pq}=\frac{1-\beta}{p}\in(0,\frac{1}{p})$. Hence it is enough to show that
$\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\frac{\beta q}{p}}\leq C(p,q).$
Step 2. Since $\beta q>2$, $\frac{\beta q}{p}>1$. By the convexity of the $\ell^{\frac{\beta q}{p}}$ with $\frac{\beta q}{p}>1$, we split the sum
$\begin{split}&\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\frac{\beta q}{p}}\\ &\leq C\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}\cap\{|f|>|\mathcal{C}_{k}^{j}|^{-1/2}\}}|f|^{p}\right)^{\frac{\beta q}{p}}\\ &+C\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}\cap\{|f|\leq|\mathcal{C}_{k}^{j}|^{-1/2}\}}|f|^{p}\right)^{\frac{\beta q}{p}}\\ &:=C(A+B),\end{split}$
where $C=C(p,q)$. We estimate the first term. For $j\in\mathbb{Z}$, define
$f^{j}=f1_{\{|f|>|\mathcal{C}_{k}^{j}|^{-1/2}\}}.$
Then
$A=\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}\frac{\beta q}{p}}\left(\int_{\mathcal{C}_{k}^{j}}|f^{j}|^{p}\right)^{\frac{\beta q}{p}}.$
Then by the comparison of the $\ell^{\frac{q\beta}{p}}$ norm and the Fubini theorem, we have
$\begin{split}A&\leq\left(\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\int_{\mathcal{C}_{k}^{j}}|f^{j}|^{p}\right)^{\frac{\beta q}{p}}\\ &=\left(\sum_{j}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\int|f^{j}|^{p}\right)^{\frac{\beta q}{p}}\\ &\leq\left(\int_{S^{d}}|f|^{p}\sum_{j:\,|f|>|\mathcal{C}_{k}^{j}|^{-1/2}}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\right)^{\frac{\beta q}{p}}\end{split}$
Since 2p>02-p>0, we can sum the geometric series and obtain
$A\leq C\left(\int_{S^{d}}|f|^{p}|f|^{2-p}\right)^{\frac{\beta q}{p}}\leq C\left(\int_{S^{d}}|f|^{2}\right)^{\frac{\beta q}{p}}\leq C$
because $\|f\|_{L^{2}(\Gamma,\sigma)}=1$. We now estimate the term $B$. For $j\in\mathbb{Z}$, define $f_{j}=f1_{|f|\leq|\mathcal{C}_{k}^{j}|^{-1/2}}$. Then
$B=\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\left(\int_{\mathcal{C}_{k}^{j}}|f_{j}|^{p}\right)^{\frac{\beta q}{p}}.$
We use the Hölder inequality with exponents $\frac{\beta q}{p}$ and $\frac{\beta q}{\beta q-p}$ to the term $\int_{\mathcal{C}_{k}^{j}}|f_{j}|^{p}$:
$\begin{split}B&\leq\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2p}\beta q}\int_{\mathcal{C}_{k}^{j}}|f_{j}|^{\beta q}\left(|\mathcal{C}_{k}^{j}|^{\frac{\beta q-p}{\beta q}}\right)^{\frac{\beta q}{p}}\\ &=\sum_{j,k}|\mathcal{C}_{k}^{j}|^{\frac{\beta q}{2}-1}\int_{\mathcal{C}_{k}^{j}}|f_{j}|^{\beta q}\\ &=\sum_{j,k}2^{jd(1-\frac{\beta q}{2})}\int_{\mathcal{C}_{k}^{j}}|f_{j}|^{\beta q}\leq\sum_{j}2^{jd(1-\frac{\beta q}{2})}\int_{S^{d}}|f_{j}|^{\beta q}\\ &=\int_{S^{d}}|f|^{\beta q}\sum_{j:\,|f|\leq 2^{jd/2}}2^{jd(1-\frac{\beta q}{2})}\leq\int_{S^{d}}|f|^{2}\leq 1,\end{split}$
since $1-\frac{\beta q}{2}<0$ and we are summing a geometric series. Then Lemma 2.4 follows. ∎
3. Profile decomposition: the first decompostion
We first state several definitions as in [9].
Definition 3.1.
Let $\mathcal{C}=\mathcal{C}(z,r)$ be a cap. For $z\in S^{d}$, define
$\Psi_{z}(x)=r^{-1}L_{z}(\Pi_{H_{z}}(x))$
for all $x$ in the hemisphere $\{x:\,x\cdot z\geq 0\}$, where $\Pi_{H_{z}}$ is the orthogonal projection onto $H_{z}$, and $L_{z}:H_{z}\to\mathbb{R}^{2}$ is an arbitrary linear isometry. The rescaled map associated with $\mathcal{C}$ is defined. by $\Phi_{\mathcal{C}_{z,r}}=\Psi_{z}^{-1}$.
Remark 3.2.
The map $\Phi_{\mathcal{C}(z,r)}:\,B(0,r^{-1})\to\text{`` the indicated hemisphere"}$. For instance, for $z=(0,0,1)$,
$\Phi_{\mathcal{C}(z,r)}(y_{1},y_{2})=(ry_{1},ry_{2},(1-r^{2}|y|^{2})^{1/2}),\forall y\in B(0,r^{-1}).$
Definition 3.3.
Let $\mathcal{C}=\mathcal{C}(z,r)$ be a cap. For $f\in L^{2}(S^{d})$, define the pullback of $f$ by
$\Phi_{\mathcal{C}}^{*}f(y)=r^{d/2}\bigl(f\circ\Phi_{\mathcal{C}}\bigr)(y).$
Remark 3.4.
We compute the following $L^{2}$ norms and find that $\|\Phi_{\mathcal{C}}^{*}f\|^{2}_{L^{2}(\mathbb{R}^{d})}$ and $\|f\|_{L^{2}(\Gamma,\sigma)}$ are comparable.
[…]
Here $J_{y}=\sqrt{1-r^{2}|y|^{2}}$. This formula and the last two lines in (11) use [31, Eq. (68) on page 498]. We see that if $r<1/2$, then $\frac{\sqrt{3}}{2}\leq J_{y}\leq 1$.
As a consequence of the refinement in Lemma 2.2, we have
Proposition 3.5.
For any $\delta>0$ there exists $C_{\delta}<\infty$ and $\eta_{\delta}>0$ with the following properties. If $f_{\nu}\in L^{2}(S^{d})$ satisfies $\|\widehat{f_{\nu}\sigma}\|_{2+4/d}\geq\delta\mathcal{R}\|f_{\nu}\|_{2}$, then for any $\delta>0$, there exists $\{f^{j}_{\nu}\}$, $\mathcal{C}_{\nu}^{j}$ and $e_{\nu}$ satisfying that
(12) $f_{\nu}=\sum_{j=1}^{N}f_{\nu}^{j}+e_{\nu}^{N},$
(13) $\limsup_{\nu}\frac{r_{\nu}^{j}}{r_{\nu}^{k}}+\frac{r_{\nu}^{k}}{r_{\nu}^{j}}+\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}=\infty,\,\|f_{\nu}\|_{2}^{2}=\sum_{j=1}^{N}\|f_{\nu}^{j}\|_{2}^{2}+\|e_{\nu}^{N}\|_{2}^{2},$
(14) $|f_{\nu}^{j}|\leq C_{\delta}\|f_{\nu}\|_{2}|\mathcal{C}_{\nu}^{j}|^{-1/2}\chi_{\mathcal{C}_{\nu}^{j}},\,\|f_{\nu}^{j}\|_{2}\geq\eta_{\delta}\|f_{\nu}\|_{2},$
(15) $\|\widehat{e^{N}\sigma}\|_{2+4/d}\leq\delta\mathcal{R}\|f_{\nu}\|_{2}.$
Here both $C^{-1}_{\delta}$ and $\eta_{\delta}$ are proportional to $\delta^{O(1)}$.
The proof uses the refined Strichartz estimate, Lemma 2.2 and iterations. During the iterations, the orthogonality relation is extracted along subsequences.
Proof.
Step 1. Let $f_{\nu}$ be such that $\|f_{\nu}\|_{L^{2}(\Gamma,\sigma)}=1$ and $0<\delta<1$. Then from Lemma 2.2, we see that
$\left(\sup_{\mathcal{C}_{\nu}}|\mathcal{C}_{\nu}|^{-1/2}\int_{\mathcal{C}_{\nu}}|f_{\nu}|d\sigma\right)^{\alpha}\geq c\delta,$
for some absolute constant $0<c$. That is to say,
$\sup_{\mathcal{C}_{\nu}}|\mathcal{C}_{\nu}|^{-1/2}\int_{\mathcal{C}_{\nu}}|f_{\nu}|d\sigma\geq c\delta^{1/\alpha}.$
Fix a cap $\mathcal{C}_{\nu}$ such that
$|\mathcal{C}_{\nu}|^{-1/2}\int_{\mathcal{C}_{\nu}}|f_{\nu}|d\sigma\geq c_{0}\delta^{1/\alpha},$
for some constant $c_{0}=\frac{c}{2}$. That is to say,
$\int_{\mathcal{C}_{\nu}}|f_{\nu}|d\sigma\geq c_{0}\delta^{1/\alpha}|\mathcal{C}_{\nu}|^{1/2}.$
Let $R_{\nu}\geq 1$. Define $E_{\nu}=\{x\in\mathcal{C}_{\nu}:|f_{\nu}(x)|\leq R\}$. Set $g_{\nu}=f_{\nu}1_{E_{\nu}}$ and $h_{\nu}=f_{\nu}-f_{\nu}1_{E_{\nu}}$. Then $g_{\nu}$ and $h_{\nu}$ have disjoint supports, $g_{\nu}+h_{\nu}=f_{\nu}$, $g_{\nu}$ is supported on $\mathcal{C}_{\nu}$, and $\|g_{\nu}\|_{\infty}\leq R$. Now $|h_{\nu}(x)|\geq R_{\nu}$ for almost every $x\in\mathcal{C}_{\nu}$ for which $h_{\nu}(x)\neq 0$, so
$\int_{\mathcal{C}_{\nu}}|h_{\nu}|\leq R^{-1}\int_{\mathcal{C}_{\nu}}|h_{\nu}|^{2}\leq{R_{\nu}}^{-1}\|f_{\nu}\|_{2}^{2}=R_{\nu}^{-1}.$
Define $R_{\nu}$ by $R_{\nu}^{-1}=\frac{c_{0}}{2}\delta^{1/\alpha}|\mathcal{C}_{\nu}|^{1/2}$. Then
$\int_{\mathcal{C}_{\nu}}|g_{\nu}|=\int_{\mathcal{C}_{\nu}}|f_{\nu}|-\int_{\mathcal{C}_{\nu}}|h_{\nu}|\geq\frac{c_{0}}{2}\delta^{1/\alpha}|\mathcal{C}_{\nu}|^{1/2}.$
By Hölder’s inequality, since $g_{\nu}$ is supported on $\mathcal{C}_{\nu}$, we have
$\|g_{\nu}\|_{2}\geq|\mathcal{C}_{\nu}|^{-1/2}\|g_{\nu}\|_{L^{1}(\mathcal{C}_{\nu})}\geq\frac{c_{0}}{2}\delta^{1/\alpha}=\frac{c_{0}}{2}\delta^{a/\alpha}\|f_{\nu}\|_{L^{2}}.$
Then the decomposition $f_{\nu}=g_{\nu}+h_{\nu}$ satisfies the conclusion that, for every $\delta>0$, there exists $C_{\delta}>0$ and $\eta_{\delta}>0$, such that for every $\delta$-nearly extremizer, there exists caps $\mathcal{C}_{\nu}$ such that
(16) $\begin{split}&0\leq|g_{\nu}|,|h_{\nu}|\leq|f_{\nu}|,\\ &\text{ the functions }g,h\text{ have disjoint supports},\\ &|g_{\nu}(x)|\leq C_{\delta}\|f_{\nu}\|_{L^{2}}|\mathcal{C}_{\nu}|^{-1/2}1_{\mathcal{C}_{\nu}}(x),\\ &\|g_{\nu}\|_{L^{2}}\geq\eta_{\delta}\|f_{\nu}\|_{L^{2}}.\end{split}$
Step 2. It is easy to see that
(17) $\Phi_{\mathcal{C}_{\nu}}^{*}(g_{\nu})(y)=r_{\nu}^{d/2}\left(g_{\nu}\circ\Phi_{\mathcal{C}_{\nu}}\right)(y)\leq C_{\delta}1_{B(0,1)}(y),\forall y\in\mathbb{R}^{d}.$
Also we have
(18) $\|f_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}=\|f_{\nu}-g_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}+\|g_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}$
since the supports are disjoint from the sphere $S^{d}$. We repeat the same arguments with $f_{\nu}-g_{\nu}$ in place of $g_{\nu}$. At each step, the $L^{2}$-norm decreases of at least $c\delta^{1/\alpha}$. With the same constant $c$ as for the first step. After $N(\delta)$ steps, we obtain $(g_{\nu}^{j})_{1\leq j\leq N(\delta)}$ and $(\gamma_{\nu}^{j})_{1\leq j\leq N(\delta)}=(z_{\nu}^{j},r_{\nu}^{j})$ satisfying the relation (17)
(19) $\begin{split}f_{\nu}&=\sum_{j=1}^{N(\delta)}g_{\nu}^{j}+q_{\nu},\\ \|\widehat{q_{\nu}\sigma}\|_{L^{2+\frac{4}{d}}(\mathbb{R}^{d+1})}&\leq\delta,\\ \|f_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}&=\sum_{j=1}^{N(\delta)}\|g^{j}_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}+\|q_{\nu}\|^{2}_{L^{2}(\Gamma,\sigma)}.\end{split}$
To see the orthogonality of parameters, we reorganize the decomposition. We say that $\gamma_{\nu}^{j}$ and $\gamma_{\nu}^{k}$ are orthogonal if for $j\neq k$, we have
(20) $\frac{r_{\nu}^{j}}{r_{\nu}^{k}}+\frac{r_{\nu}^{k}}{r_{\nu}^{j}}+\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}\to\infty,\text{ as }\nu\to\infty.$
Define
$h_{\nu}^{1}=\sum_{j=1}^{N(\delta)}g_{\nu}^{j}-\sum_{\gamma_{\nu}^{j}\perp\gamma_{\nu}^{1}}g_{\nu}^{j}.$
If there exists $2\leq j_{0}\leq N(\delta)$ such that $\gamma_{\nu}^{j_{0}}\perp\gamma_{\nu}^{1}$, then we define
$h_{\nu}^{2}=\sum_{j=1}^{N(\delta)}g_{\nu}^{j}-\sum_{\gamma_{\nu}^{j}\perp\gamma_{\nu}^{1}\atop\gamma_{\nu}^{j}\perp\gamma_{\nu}^{j_{0}}}g_{\nu}^{j}.$
Repeating this argument a finite number of times, we rearrange the above sum. The $L^{2}$-norm orthogonality relation in (19) holds, since the supports of the functions we consider are disjoint on the sphere. The $g_{\nu}^{j}$’s kept in the definition $h_{\nu}^{1}$ are such that $r_{\nu}^{j}$ are not orthogonal one to another. It is sufficiently to show that up to a subsequence, $\Phi_{\mathcal{C}_{\nu}^{1}}(g_{\nu}^{j})$ is bounded. By construction, $\Phi_{\mathcal{C}_{\nu}^{j}}^{j*}(g_{\nu}^{j})$ are bounded by $c_{\delta}1_{B(0,1)}(y)$.
We compute, $\forall y\in S^{d}$ and $f$ supported on the sphere,
(21) $\begin{split}\Phi_{\mathcal{C}_{\nu}^{1}}^{*}\left(\bigl(\Phi_{\mathcal{C}_{\nu}^{j}}^{*}\bigr)^{-1}f\right)(y)&=(r_{\nu}^{1})^{d/2}\left[\bigl(\bigl(\Phi_{\mathcal{C}_{\nu}^{j}}^{*}\bigr)^{-1}f\bigr)\circ\Phi_{\mathcal{C}_{\nu}^{1}}\right](y)\\ &=(r_{\nu}^{1})^{d/2}\bigl(\bigl(\Phi_{\mathcal{C}_{\nu}^{j}}^{*}\bigr)^{-1}f\bigr)\bigl(\Pi_{H_{z_{\nu}^{1}}}^{-1}\circ L_{1}^{-1}(r_{\nu}^{1}y)\bigr)\\ &=\bigl(\frac{r_{\nu}^{1}}{r_{\nu}^{j}}\bigr)^{d/2}f\left((r_{\nu}^{1})^{-1}L_{j}\left(\Pi_{z_{\nu}^{j}}\bigl(\Pi_{H_{z_{\nu}^{1}}}^{-1}\circ L_{1}^{-1}(r_{\nu}^{1}y)\bigr)\right)\right).\end{split}$
This is the Euclidean coordinate representation of the function $f$ on the sphere $S^{d}$. Now we show that it is bounded by $C_{\delta}1_{B(0,1)}(y).$ If
$\sup_{\nu}\left\{\frac{r_{\nu}^{1}}{r_{\nu}^{j}}+\frac{r_{\nu}^{j}}{r_{\nu}^{1}}+\frac{|z_{\nu}^{1}-z_{\nu}^{j}|}{r_{\nu}^{j}}\right\}<\infty,$
then up to a subsequence, we may assume that
$\frac{r_{\nu}^{1}}{r_{\nu}^{j}}\to c_{1},$
for some $0<c_{1}<\infty$. The same applies to $\frac{|z_{\nu}^{1}-z_{\nu}^{j}|}{r_{\nu}^{j}}$, that is to say, up to a subsequence,
$\frac{|z_{\nu}^{1}-z_{\nu}^{j}|}{r_{\nu}^{j}}\to c_{2}$
for some $c_{2}\geq 0$. Then
$r_{\nu}^{1}\leq Cr_{\nu}^{j},|z_{\nu}^{1}-z_{\nu}^{j}|<Cr_{\nu}^{j}.$
If $z\in\mathcal{C}(z_{\nu}^{1},r_{\nu}^{1})$, then
$|z-z_{\nu}^{j}|\leq Cr_{\nu}^{j}.$
Therefore the term in (21) is less than or equal to
$C_{\delta}1_{B(0,1)}(y).$
The same applies to $\Phi_{\mathcal{C}_{\nu}^{1}}^{j*}(g_{\nu}^{k})$ for the terms $k$ such that $\gamma_{\nu}^{k}\not\perp\gamma_{\nu}^{1}.$ Therefore the proof for the Proposition 3.5 is complete. ∎
4. The second decomposition for each piece
For each component $f_{\nu}^{j}$ above, we decompose it further and establish the orthogonality of profiles.
4.1. Key propositions
In this section, we may assume that the function $f_{\nu}^{j}$ is supported on a cap $\mathcal{C}_{\nu}^{j}(z_{\nu}^{j},r_{\nu}^{j})$. The decomposition for $\widehat{f_{\nu}^{j}\sigma}$ is motivated by the rescaling relation,
(22) $\begin{split}&\left|\widehat{f_{\nu}^{j}\sigma}(x,t)\right|=\left|\int_{\mathcal{C}_{\nu}^{j}}e^{i(x,t)\cdot\xi}f_{\nu}^{j}(\xi)d\sigma(\xi)\right|\\ &=(r_{\nu}^{j})^{d/2}\left|\int_{\mathbb{R}^{d}}e^{ix\cdot r_{\nu}^{j}y+it\sqrt{1-|r_{\nu}^{j}y|^{2}}}\frac{(r_{\nu}^{j})^{d/2}f_{\nu}^{j}\circ\Pi^{-1}_{H_{z_{\nu}^{j}}}\circ L^{-1}_{z_{\nu}^{j}}(r_{\nu}^{j}y)}{\sqrt{1-|r_{\nu}^{j}y|^{2}}}dy\right|\\ &=\left|(r_{\nu}^{j})^{d/2}\int e^{i(r_{\nu}^{j})x\cdot y-\frac{it(r_{\nu}^{j})^{2}|y|^{2}}{2}}e^{i(r_{\nu}^{j})^{2}t\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{1}{(1-(r_{\nu}^{j})^{2}|y|^{2})^{1/4}}\right.\times\\ &\left.\qquad\qquad\qquad\qquad\qquad\times\dfrac{(r_{\nu}^{j})^{d/2}f_{\nu}^{j}\circ\Pi^{-1}_{H_{z_{\nu}^{j}}}\circ L^{-1}_{z_{\nu}^{j}}(r_{\nu}^{j}y)}{(1-(r_{\nu}^{j})^{2}|y|^{2})^{1/4}}dy\right|\\ &=\left|(r_{\nu}^{j})^{d/2}e^{\frac{i(r_{\nu}^{j})^{2}t\Delta}{2}}\bigl(h_{\nu}((r_{\nu}^{j})^{2}t,\cdot)g^{j}_{\nu}(\cdot)\bigr)(r_{\nu}x)\right|,\end{split}$
where
(23) $\begin{split}h_{\nu}^{j}(t,y)&:=e^{it\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{1}{(1-(r_{\nu}^{j})^{2}|y|^{2})^{1/4}},\\ g_{\nu}^{j}(y)&:=\dfrac{(r_{\nu}^{j})^{d/2}f_{\nu}^{j}\circ\Pi^{-1}_{H_{z_{\nu}^{j}}}\circ L^{-1}_{z_{\nu}^{j}}(r_{\nu}^{j}y)}{(1-(r_{\nu}^{j})^{2}|y|^{2})^{1/4}}.\end{split}$
So a decomposition for $\widehat{f_{\nu}^{j}\sigma}$ immediately follows once we have a decomposition for $\{g_{\nu}^{j}\}$. The definitions in (23) is reminiscent of [14, Equation (2.3)]. In what follows in this section, we will suppress the superscripts $j$.
Proposition 4.2.
Let $\{g_{\nu}\}$ be defined as above and let $C>0$ be the universal constant such that
(24) $|g_{\nu}|\leq C1_{B(0,1)}.$
Then there exists a sequence of functions $\{\phi^{j}\}_{j}\in L^{2}$ and $(x_{\nu}^{j},t_{\nu}^{j})\in{\mathbb{R}}^{d}\times\mathbb{R}$ and $e_{\nu}^{l}\in L^{2}$ such that
(25) $g_{\nu}(y)=\sum_{j=1}^{l}e^{\frac{it^{j}_{\nu}|y|^{2}}{2}}e^{-ix^{j}_{\nu}y}\phi^{j}+e^{l}_{\nu}(y).$
with the following properties: $\phi^{j}$ can be taken to be smooth functions supported on a ball of radius 1 uniformly in $\nu$. The parameters $\{(x_{\nu}^{j},t_{\nu}^{j})\}$ satisfy, for $k\neq j$,
(26) $|x_{\nu}^{k}-x_{\nu}^{j}|+|t_{\nu}^{k}-t_{\nu}^{j}|\to\infty,\text{ as }\nu\to\infty.$
For each $l\geq 0$,
(27) $\limsup_{\nu\to\infty}\left(\|g_{\nu}\|_{L^{2}(\Gamma,\sigma)}^{2}-\sum_{j=1}^{l}\|\phi^{j}\|_{L^{2}(\mathbb{R}^{d})}^{2}-\|e_{\nu}^{l}\|^{2}_{L^{2}}\right)=0.$
where $e_{\nu}^{l}$ satisfies
(28) $\limsup_{l\to\infty}\limsup_{\nu\to\infty}\left\|r_{\nu}^{d/2}e^{\frac{itr_{\nu}^{2}\Delta}{2}}\bigl(h(r_{\nu}^{2}t,\cdot)e_{\nu}^{l}\bigr)(r_{\nu}x)\right\|_{L^{2+4/d}(\mathbb{R}^{d+1})}=0.$
where
$e^{\frac{it\Delta}{2}}f(x)=\int e^{ixy-\frac{it|y|^{2}}{2}}f(y)dy.$
The following result is for two profiles with the same center $z_{\nu}^{j}$ and $r_{\nu}^{j}$. More general decomposition is obtained in the next section when we combine Propositions 3.5 and 4.2.
Proposition 4.3 (Orthogonality I).
Let $\{h_{\nu}\}$ be defined as above and set
(29) $\begin{split}G_{\nu}^{k}(y)&:=e^{\frac{it^{k}_{\nu}|y|^{2}}{2}}e^{-ix^{k}_{\nu}y}\phi^{k},\\ G_{\nu}^{j}(y)&:=e^{\frac{it^{j}_{\nu}|y|^{2}}{2}}e^{-ix^{j}_{\nu}y}\phi^{j}.\end{split}$
Then for $k\neq j$,
(30) $\begin{split}&\limsup_{\nu\to\infty}\left\|\Bigl(r_{\nu}^{d/2}e^{\frac{itr_{\nu}^{2}\Delta}{2}}\bigl(h_{\nu}(r_{\nu}^{2}t,\cdot)G_{\nu}^{k}\bigr)(r_{\nu}x)\Bigr)\Bigl(r_{\nu}^{d/2}e^{\frac{itr_{\nu}^{2}\Delta}{2}}\bigl(h_{\nu}(r_{\nu}^{2}t,y)G_{\nu}^{j}\bigr)(r_{\nu}x)\Bigr)\right\|_{L_{t,x}^{1+2/d}(\mathbb{R}^{d})}=0,\end{split}$
where the integrand $\bigl(r_{\nu}^{d/2}e^{\frac{itr_{\nu}^{2}\Delta}{2}}\bigl(h_{\nu}(r_{\nu}^{2}t,\cdot)G_{\nu}^{k}\bigr)(r_{\nu}x)$ is understood as the Fourier transform of a surface measure on the sphere.
Proof of Proposition 4.2..
We split the proof into two steps. We will follow the proof in [5, Proposition 3.4] and present them for sake of completeness.
Step 1. For $(x_{\nu},t_{\nu})\in{\mathbb{R}}^{d+1}$, we define
$T_{\nu}(g)(y)=e^{-\frac{it_{\nu}|y|^{2}}{2}}e^{ix_{\nu}y}g(y);$
analogously $T_{\nu}^{i}$ for $(x_{\nu}^{i},t_{\nu}^{i})$ for $i\geq 1$, and $T_{\nu}^{-1}(g)(y)=e^{\frac{it_{\nu}|y|^{2}}{2}}e^{-ix_{\nu}y}g(y)$. Let $P^{0}$ denote the sequence $\{g_{\nu}\}_{\nu\geq 1}$. Then we define the set
$\mathcal{W}(P^{0})=\{w-\lim_{\nu\to\infty}T_{\nu}(P_{\nu}^{0})(y)\text{ in }L^{2}:\,(x_{\nu},t_{\nu})\in{\mathbb{R}}^{d+1}\},$
where $w-\lim g_{\nu}$ denotes a weak limit of $\{g_{\nu}\}$ in $L^{2}$. Define the blow-up criterion associated to $\mathcal{W}(P^{0})$:
$\mu(P^{0}):=\sup\{\|\phi\|_{L^{2}}:\phi\in\mathcal{W}(P^{0})\}.$
Then for any $\phi\in\mathcal{W}(P^{0})$, $\|\phi\|_{L^{2}}\leq\limsup\|g_{\nu}\|_{L^{2}(\mathbb{R}^{d})}=\limsup\|f_{\nu}\|_{L^{2}(\sigma)}$, which folllows from [18, Theorem 2.11] about the $L^{2}$-norm comparison for a weak convergent sequence.
If $\mu(P^{0})=0$, then we set $l=0$, and $e_{\nu}^{0}=g_{\nu}$ for all $\nu\geq 1$. Otherwise, $\mu(P^{0})>0$, then up to a subsequence, there exists nontrivial $\phi^{1}\in L^{2}$ and $(x_{\nu}^{1},t_{\nu}^{1})_{\nu\geq 1}$ such that
(31) $\phi^{1}$ $=w-\lim_{\nu\to\infty}T_{\nu}^{1}(P_{\nu}^{0})(y),$
(32) $\|\phi^{1}\|_{2}$ $\geq\frac{1}{2}\mu(P^{0}).$
Let $P^{1}$ denote the sequence $\{g_{\nu}(y)-(T_{\nu}^{1})^{-1}(\phi^{1})(y)\}_{\nu\geq 1}$. It is not hard to see that
(33) $w-\lim_{\nu\to\infty}T_{\nu}^{1}(P_{\nu}^{1})=0,$
(34) $\|g_{\nu}\|_{L^{2}(\sigma)}^{2}-\|\phi^{1}\|_{2}^{2}-\|e_{\nu}^{1}\|^{2}_{L^{2}}\to 0,\text{ as }\nu\to\infty.$
Indeed, (33) follows from the definition. For (34),
$\|e_{\nu}^{1}\|_{L^{2}}^{2}=\|T_{v}^{1}(g_{\nu})-\phi^{1}\|_{L^{2}}^{2}=\|T_{\nu}^{1}(g_{\nu})-\phi^{1}\|_{L^{2}}^{2}-2\operatorname{Re}\langle T_{\nu}^{1}(g_{\nu}),\phi^{1}\rangle+\|\phi^{1}\|_{L^{2}}^{2}.$
It implies that
$\|g_{\nu}\|_{L^{2}}^{2}-2\operatorname{Re}\langle T_{\nu}^{1}(g_{\nu}),\phi^{1}\rangle+\|\phi^{1}\|_{L^{2}}^{2}-\|e_{\nu}^{1}\|_{L^{2}}^{2}\to\|g_{\nu}\|_{L^{2}(\sigma)}^{2}-\|\phi^{1}\|_{2}^{2}-\|e_{\nu}^{1}\|^{2}_{L^{2}}\to 0,\text{ as }\nu\to\infty.$
For $P^{1}=\{g_{\nu}(y)-(T_{\nu}^{1})^{-1}(\phi^{1})(y)\}_{\nu\geq 1}$, we iteratively consider the set
$\mathcal{W}(P^{1})=\{w-\lim_{\nu\to\infty}T_{\nu}(P_{\nu}^{1})\text{ in }L^{2}:\,(x_{\nu},t_{\nu})\in{\mathbb{R}}^{d+1}\}.$
Then we test whether $\mu(P^{1})>0$: if $\mu(P^{1})=0$, then the algorithm stops; take $l=1$, and $e_{\nu}^{1}=P^{1}_{\nu}$; if not, then up to a subsequence, there exists nontrivial $\phi^{2}\in L^{2}$ and and $(x_{\nu}^{2},t_{\nu}^{2})_{\nu\geq 1}$ such that
(35) $\phi^{2}$ $=w-\lim_{\nu\to\infty}T_{\nu}^{2}(P_{\nu}^{1})(y),$
(36) $\|\phi^{2}\|_{2}$ $\geq\frac{1}{2}\mu(P^{1}).$
By similar considerations as in (33) and (34), if setting $P_{\nu}^{2}=P_{\nu}^{1}-(T_{\nu}^{2})^{-1}(\phi^{2})$, then
$w-\lim_{\nu\to\infty}T_{\nu}^{2}(P_{\nu}^{2})=0,$
$\|g_{\nu}\|_{L^{2}(\sigma)}^{2}-\sum_{j=1}^{2}\|\phi^{j}\|_{2}^{2}-\|e_{\nu}^{2}\|_{L^{2}}^{2}\to 0,\text{ as }\nu\to\infty.$
Now we claim that (26) holds. Otherwise, up to a subsequence we may assume that
$|t_{\nu}^{2}-t_{\nu}^{1}|+|x_{\nu}^{2}-x_{\nu}^{1}|\to c,\text{ as }\nu\to\infty,$
for some $0<c<\infty$. In this case, the dominated convergence theorem gives, up to a subsequence,
$T_{\nu}^{2}(T_{\nu}^{1})^{-1}\text{ converges strongly in }L^{2}.$
This will imply that
(37) $T_{\nu}^{2}(P_{\nu}^{1})\to 0,\text{ weakly in }L^{2},$
as $T_{\nu}^{1}(P_{\nu}^{1})\to 0$ weakly in $L^{2}$ and the following relation holds,
$T_{\nu}^{2}(P_{\nu}^{1})=T_{\nu}^{2}(T_{\nu}^{1})^{-1}\bigl(T_{\nu}^{1}(P_{\nu}^{1})\bigr).$
But the claim in (37) is a contradiction to the existence of nontrivial $\phi^{2}$. So (26) holds.
Iterating this process, a diagonalization argument produces a family of pairwise orthogonal sequences $(x_{\nu}^{j},t_{\nu}^{j})$ and $\phi^{j}$ satisfying (25), (26) and (27). Since $\sum_{j}\|\phi^{j}\|_{2}^{2}\leq\sup_{\nu}\|g_{\nu}\|_{2}^{2}<\infty$ and $\mu(P^{l+1})\leq 2\|\phi^{l}\|_{2}$, we have
(38) $\mu(P^{l})\to 0,\text{ as }l\to\infty.$
We give the following two remarks. (1). Since $g_{\nu}$ is supported on the unit ball and bounded, $\phi^{j}$ can be taken to be smooth functions supported on a ball of radius 1 uniformly in $\nu$. (2). The functions $\{e_{\nu}^{l}\}$ is supported in the unit ball and uniformly bounded in $l$ and $\nu$. Note that the proof of (27) yields a more precise result than (27): for $l\geq 1$ and for $\psi\in L^{2}\cap L^{\infty}(\mathbb{R}^{d})$,
(39) $\limsup_{\nu\to\infty}\left(\|\psi g_{\nu}\|_{L^{2}(\Gamma,\sigma)}^{2}-\sum_{j=1}^{l}\|\psi\phi^{j}\|_{L^{2}(\mathbb{R}^{d})}^{2}-\|\psi e_{\nu}^{l}\|^{2}_{L^{2}}\right)=0.$
Then for the support of $\{e_{\nu}^{l}\}$, we consider $1_{E}$ for $E\subset\bigl(B(0,1)\bigr)^{c}$ and the measure of $E$, $|E|>0$, the Lebesgue measure of $E$. Then because $g_{\nu}$ and $\phi^{j}$ are supported in $B(0,1)$,
$\|1_{E}e_{\nu}^{l}\|_{L^{2}}=0,\text{ as }\nu\to\infty,$
which implies
[…]
Secondly for the boundedness of $\{e_{\nu}^{l}\}$: given any $A>0$, let
$E_{A}:=\{x\in B(0,1):\,|e_{\nu}^{l}|>A\}.$
Then we have
$A^{2}|E_{A}|\leq\|1_{E_{A}}e_{\nu}^{l}\|_{L^{2}}^{2}\leq\|1_{E_{A}}g_{\nu}\|_{L^{2}}^{2}\leq\|g_{\nu}\|_{L^{\infty}}^{2}|E_{A}|.$
If $|E_{A}|\neq 0$, then $A\leq\|g_{\nu}\|_{L^{\infty}}.$ This proves that $\|e_{\nu}^{l}\|_{L^{\infty}}\leq\|g_{\nu}\|_{L^{\infty}}$ for sufficiently large $\nu$; for the definition of the $L^{\infty}$ norm, see e.g.,[12, Page 184]. Then a diagonalization argument yields that the functions $\{e_{\nu}^{l}\}$ is supported in the unit ball and uniformly bounded in $l$ and $\nu$.
Step 2. At this step, we show that some localized restriction estimate $L^{\infty}(S^{d})\to L^{q}_{t,x}$ for some $q<2+4/d$ and the information that $\lim\mu(P^{l})=0$ will imply (28). The restriction estimate we need is
Theorem F.
There exists $q<2+4/d$. If $f\in L^{\infty}(S^{d})$ for a large $R>0$, then
$\|\widehat{f\sigma}\|_{L^{q}_{t,x}}\leq C\|f\|_{L^{\infty}(\sigma)}.$
Similarly if $f\in L^{\infty}(B(0,R))$, then
$\|e^{it\Delta}f\|_{L^{q}_{t,x}}\leq C_{R}\|f\|_{L^{\infty}}.$
This follows from the adjoint linear restriction estimate for the sphere beyond Tomas-Stein’s inequality in [33].
We continue proving (28). There are still two cases to consider: $r_{\nu}\to 0$ or $r_{\nu}\to r_{0}>0$. It suffices to consider the case $r_{\nu}\to 0$ as the other case is similar. Then by scaling, the norm on the left hand side of (28) is equivalent to
$\|e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t)e_{\nu}^{l}\bigr)\|_{L^{2+4/d}_{t,x}}.$
From the way $e_{\nu}^{l}$ is obtained, we may assume that $e_{\nu}^{l}$ is compactly supported in the unit ball, and bounded uniformly in $\nu$ and $l$. Then by Lemma F and the real interpolation [30, Theorem 1.3], (28) is reduced to
(40) $\limsup_{l\to\infty}\limsup_{\nu\to\infty}\|e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t)e_{\nu}^{l}\bigr)\|_{L^{\infty}_{t,x}}=0.$
This follows from the fact that $\mu(P^{l})\to 0$ as $l\to\infty.$ Indeed, there exists $(x^{l}_{\nu},t^{l}_{\nu})$ such that, up to a subsequence,
(41) $\left|e^{\frac{it_{\nu}^{l}\Delta}{2}}\bigl(h_{\nu}(t_{\nu}^{l})e_{\nu}^{l}\bigr)(x_{\nu}^{l})\right|\sim\|e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t)e_{\nu}^{l}\bigr)\|_{L^{\infty}_{t,x}}.$
On the other hand, since $e_{\nu}^{l}$ is compactly supported,
$e^{-\frac{it^{l}_{\nu}|y|^{2}}{2}}e^{ix^{l}_{\nu}y}h_{\nu}(t_{\nu}^{l})e_{\nu}^{l}(y)=e^{-\frac{it^{l}_{\nu}|y|^{2}}{2}}e^{ix^{l}_{\nu}y}h_{\nu}(t_{\nu}^{l})e_{\nu}^{l}(y)\phi_{R}(y)$
for some suitable bump function $\phi$ adapted to the ball $B(0,1)$; taking integration in $y$ on both sides, we have
(42) $e^{\frac{it_{\nu}^{l}\Delta}{2}}\bigl(h_{\nu}(t_{\nu}^{l})e_{\nu}^{l}\bigr)(x_{\nu}^{l})=\langle e^{-\frac{it^{l}_{\nu}|y|^{2}}{2}}e^{ix^{l}_{\nu}y}h_{\nu}(t_{\nu}^{l},y)e_{\nu}^{l}(y),\,\phi\rangle_{L^{2}}.$
The right hand side of the above is equal to
$\int_{|y|\leq 1}e^{ix_{\nu}^{l}y+it_{\nu}^{l}\frac{\sqrt{1-|r_{\nu}^{l}y|^{2}}-1}{(r_{\nu}^{l})^{2}}}\frac{e_{\nu}^{l}(y)\phi(y)}{(1-|r_{\nu}^{l}y|^{2})^{1/4}}dy.$
By using the stationary phase estimate [31, Page 334, Proposition 3], since $r_{\nu}^{l}\leq\frac{1}{2}$ for all $l,\nu$, if for a $l\geq 1$, $|t_{\nu}^{l}|\to\infty$ as $\nu\to\infty$,
$\left|\int_{|y|\leq 1}e^{ix_{\nu}^{l}y+it_{\nu}^{l}\frac{\sqrt{1-|r_{\nu}^{l}y|^{2}}-1}{(r_{\nu}^{l})^{2}}}\frac{e_{\nu}^{l}(y)\phi(y)}{(1-|r_{\nu}^{l}y|^{2})^{1/4}}dy\right|\leq C|t_{\nu}^{l}|^{-d/2}\to 0,\text{ as }\nu\to\infty.$
There is also the case where $\{t_{\nu}^{l}\}_{\nu}$ is bounded, so up to a subsequence, $\{t_{\nu}^{l}\}_{\nu}$ converges to a fixed number. Then the sequence of functions $\{e_{\nu}^{l}\}_{\nu\geq 1}$ converges strongly in $L^{2}(\mathbb{R}^{d})$ to a fixed function. So if $P^{l}:=\{e_{\nu}^{l}\}_{\nu\geq 1}$, by the definition of $\mu(P^{l})$,
[…]
since $\mu(P^{l})\to 0$ as $l\to\infty.$ This finishes the proof of (28). Therefore the proof of Proposition 4.2 is complete. ∎
Next we show that (26) implies the orthogonality result (30) in Proposition 4.3.
Proof of (30) in Proposition 4.3..
For given $j,k$, we know that $\phi^{j}$ and $\phi^{k}$ are smooth functions compactly supported in some ball of radius 1.
Recall that
$e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t,y)G_{\nu}^{k}\bigr)=\int e^{i(x-x_{\nu}^{k})-\frac{i(t-t_{\nu}^{k})|y|^{2}}{2}}e^{i(t-t_{\nu}^{k})\bigl(\frac{\sqrt{1-r^{2}_{\nu}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{k}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}dy.$
Likewise for $e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t,y)G_{\nu}^{j}\bigr)$. By a change of variables, we need to show
(44) $\begin{split}\left\|e^{i\frac{t-(t_{\nu}^{j}-t_{\nu}^{k})}{2}\Delta}\Bigl(e^{i\bigl(t-(t_{\nu}^{j}-t_{\nu}^{k})\bigr)\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\bigl(x-(x_{\nu}^{j}-x_{\nu}^{k})\bigr)\right.\\ \quad\quad\times\left.e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{k}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\right\|_{L^{1+2/d}_{t,x}}\end{split}$
goes to zero as $\nu$ goes to infinity.
For a large $N\gg 1$, set
$\Omega_{N}:=\{(t,x):|t|+|x|\leq N\},\quad\Omega_{N,\nu}=\Omega_{N}-(t_{\nu}^{j}-t_{\nu}^{k},x_{\nu}^{j}-x_{\nu}^{k}).$
We first claim that, for $\Omega=\Omega_{N}$ or $\Omega_{N,\nu}$,
(45) $\begin{split}\int_{\Omega_{N}^{c}}\left|e^{i\frac{t-(t_{\nu}^{j}-t_{\nu}^{k})}{2}\Delta}\Bigl(e^{i\bigl(t-(t_{\nu}^{j}-t_{\nu}^{k})\bigr)\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\bigl(x-(x_{\nu}^{j}-x_{\nu}^{k})\bigr)\right.\\ \quad\quad\times\left.e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{k}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\right|^{1+\frac{2}{d}}dtdx\end{split}$
goes to zero as $N$ goes to infinity uniformly in $\nu$. Here $\Omega^{c}:={\mathbb{R}}^{d+1}\setminus\Omega.$
We write
(46) $\begin{split}&e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)(x)\\ &\quad=\int e^{ixy+it\frac{\sqrt{1-|r_{\nu}y|^{2}}-1}{r_{\nu}^{2}}}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}dy.\end{split}$
We observe that, for $r_{\nu}\leq\frac{1}{2}$ and all $|y|\leq 1$, $\bigl|\partial^{2}_{y_{i}}\frac{\sqrt{1-|r_{\nu}y|^{2}}-1}{r_{\nu}^{2}}\bigr|\geq c>0$ for some $c>0$ uniformly in $\nu$. Then by the stationary phase estimate [31, p.334], the quantity in (46) is bounded by $C|t|^{-d/2}$. On the other hand, by integration by parts, if $|x|>10|t|$, for all sufficiently large $\nu$,
(47) $\left|e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)(x)\right|\leq C_{N}|x|^{-N},\forall N\in\mathbb{N}.$
Furthermore there always hold a trivial bound, for all $x,t$,
(48) $\left|e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)(x)\right|\leq C.$
Here $C$ depends on the function $\phi^{j}$.
We are now ready to prove (45) when $\Omega=\Omega_{N}$; the case where $\Omega=\Omega_{N,\nu}$ is similar and so will be omitted. By Cauchy-Schwarz,
[…]
On the one hand, by Tomas-Stein’s inequality for the sphere and a change of variables the first term above equals
(50) $\begin{split}&\left\|e^{i\frac{t}{2}\Delta}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\right\|_{L^{2+4/d}_{t,x}}\\ &=\left\|\int e^{ixy+it\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}}{r_{\nu}^{2}}}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}dy\right\|_{L^{2+4/d}_{t,x}}\\ &=\left\|\int_{|y|\leq 1/2}e^{ixy+it\sqrt{1-|y|^{2}}}\bigl(r_{\nu}^{-d/2}\phi^{j}(r_{\nu}^{-1}y)(1-|y|^{2})^{1/4}\frac{dy}{\sqrt{1-|y|^{2}}}\right\|_{L^{2+4/d}_{t,x}}\\ &=\left\|\int e^{i(x,t)\xi}\bigl(r_{\nu}^{-d/2}\phi^{j}(r_{\nu}^{-1}y)\sqrt{1-|y|^{2}}\bigr)d\sigma\right\|_{L^{2+4/d}_{t,x}}\\ &\leq C\left(\int|r_{\nu}^{-d/2}\phi^{j}(r_{\nu}^{-1}y)|^{2}\sqrt{1-|y|^{2}}d\sigma(y)\right)^{1/2}\\ &\leq C\left(\int|r_{\nu}^{-d/2}\phi^{j}(r_{\nu}^{-1}y)|^{2}dy\right)^{1/2}=C\|\phi^{j}\|_{L^{2}}\end{split}$
for some $C>0$ independent of $\nu$. Then we have estimates (46), (47) and (48), which yields that
(51) $\left\|e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{k}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\right\|^{1+2/d}_{L^{2+4/d}_{t,x}(\Omega^{c}_{N})}\to 0,\text{ uniform in }\nu\text{ as }N\to\infty.$
Therefore we have established (45), which can be also established by using the Tomas-Stein inequality for the sphere and the dominated convergence theorem. To finish the proof of (44), we need to show that, for a fixed $N\gg 1$,
(52) $\begin{split}\int_{\Omega_{N}\cap\Omega_{N,\nu}}\left|e^{i\frac{t-(t_{\nu}^{j}-t_{\nu}^{k})}{2}\Delta}\Bigl(e^{i\bigl(t-(t_{\nu}^{j}-t_{\nu}^{k})\bigr)\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{j}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\bigl(x-(x_{\nu}^{j}-x_{\nu}^{k})\bigr)\right.\\ \quad\quad\times\left.e^{i\frac{t\Delta}{2}}\Bigl(e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\frac{\phi^{k}(y)}{(1-r_{\nu}^{2}|y|^{2})^{1/4}}\Bigr)\right|^{1+\frac{2}{d}}dtdx\end{split}$
goes to zero as $\nu$ goes to infinity. It actually holds as
$\operatorname{measure}(\Omega_{N}\cap\Omega_{N,\nu})\to 0,\text{ as }\lim_{\nu\to\infty}|t_{\nu}^{j}-t_{\nu}^{k}|+|x_{\nu}^{j}-x_{\nu}^{k}|=\infty,$
and we can apply $L^{\infty}_{t,x}$-bounds to both integrals. Therefore the proof of (30) is complete. ∎
5. The full decomposition
The orthogonality result of (30) in Proposition 4.3 is established under the assumption that the sequence $\{f_{\nu}^{j}\}$ is all supported on the same cap $\mathcal{C}(z_{\nu}^{j},r_{\nu}^{j})$. In this section, we combine Propositions 3.5, 4.2 to obtain a full decomposition for $\widehat{f_{\nu}\sigma}$. We need a lemma on the adjoint bilinear restriction estimate for paraboloids from Tao [33] and Bourgain [4].
Lemma 5.1.
Assume $f_{1}$ and $f_{2}$ are supported on the two caps $\mathcal{C}_{1}(z_{1},r)$ and $\mathcal{C}_{2}(z_{2},r)$ on the sphere $S^{d}$. Let $N:=\frac{|z_{1}-z_{2}|}{r}\gg 1$. Then for any $q>\frac{d+3}{d+1}$, there exists $\alpha=\alpha(d)>0$,
(53) $\|\widehat{f_{1}\sigma}\widehat{f_{2}\sigma}\|_{L^{q}_{t,x}}\lesssim\left(\frac{|z_{1}-z_{2}|}{r}\right)^{-\alpha}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}..$
Proof.
We recall the definition of the rescaled-map in Definition 3.1. We express the sphere surface measure in $\widehat{f\sigma}$ in terms of the plane coordinates, we see that
$\left|\widehat{f_{1}\sigma}(t,x)\right|$ $=\left|r^{d/2}\int e^{ix\cdot ry+ir^{2}t\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/4}}dy\right|,$
$\left|\widehat{f_{2}\sigma}(t,x)\right|$ $=\left|r^{d/2}\int e^{ix\cdot ry+ir^{2}t\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/4}}dy\right|.$
Then by the change of variables, we see that
$\|\widehat{f_{1}\sigma}\widehat{f_{2}\sigma}\|_{L^{\frac{d+2}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}$ $=\|\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy$
$\quad\quad\times\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy\|_{L^{\frac{d+2}{d}}_{t,x}}.$
By the triangle inequality $|ry_{1}-ry_{2}|\sim|z_{1}-z_{2}|$ and $y_{i}$ are points in balls of radius $1$ for $i=1,2$ with one ball centered at the origin, $|y_{1}-y_{2}|\sim\frac{|z_{1}-z_{2}|}{r}$, as indicated in Figure 2. Thus we see the supports of the two functions in the integrands are separated by $N\sim\frac{|z_{1}-z_{2}|}{r}$. By interpolation, the $L_{t,x}^{\frac{d+2}{d}}$ is controlled by the $L^{q_{1}}_{t,x}$ with some $\frac{d+3}{d+1}<q_{1}<\frac{d+2}{d}$ and $L^{2}_{t,x}$. That is to say, there exists $\theta\in(0,1)$ such that
$\frac{d}{d+2}=\frac{1-\theta}{q_{1}}+\frac{\theta}{2}.$
Figure 2. Distant caps interact weakly.
Then by Tao’s bilinear restriction estimate for elliptic phases [33, The third remark in Section 9], more precisely [33, Corollary 1.3], the $L^{q_{1}}_{t,x}$ norm
$\|\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy\|_{L^{q_{1}}_{t,x}}$
$\lesssim N^{d-\frac{d+2}{q_{1}}}\|\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}\|_{L^{2}(\mathbb{R}^{d})}\|\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}\|_{L^{2}(\mathbb{R}^{d})}$
$\lesssim N^{d-\frac{d+2}{q_{1}}}\|f_{1}\|_{L^{2}(S^{d},\sigma)}\|\|_{L^{2}(S^{d},\sigma)}.$
By the Bourgain bilinear restriction estimate [4, Lemma 111] and [16, Theorem 4.18], the $L^{2}_{t,x}$ is bounded by
$\|\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy\int e^{ix\cdot y+it\frac{\sqrt{1-|ry|^{2}}-1}{r^{2}}}\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}dy\|_{L^{2}_{t,x}}$
$\lesssim N^{-\frac{1}{2}}\|\dfrac{r^{d/2}f_{1}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}\|_{L^{2}(\mathbb{R}^{d})}\|\dfrac{r^{d/2}f_{2}\circ\Pi^{-1}_{H_{z}}\circ L^{-1}_{z}(ry)}{(1-|ry|^{2})^{1/2}}\|_{L^{2}(\mathbb{R}^{d})}$
$\lesssim N^{-\frac{1}{2}}\|f_{1}\|_{L^{2}(S^{d},\sigma)}\|\|_{L^{2}(S^{d},\sigma)}.$
Therefore
$\|\widehat{f_{1}\sigma}\widehat{f_{2}\sigma}\|_{L^{\frac{d+2}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}\lesssim\left(N^{d-\frac{d+2}{q_{1}}}\|f_{1}\|_{L^{2}(S^{d},\sigma)}\|\|_{L^{2}(S^{d},\sigma)}\right)^{1-\theta}\left(N^{d-\frac{d+2}{q_{1}}}\|f_{1}\|_{L^{2}(S^{d},\sigma)}\|\|_{L^{2}(S^{d},\sigma)}\right)^{\theta}$
$\lesssim N^{\bigl(d-\frac{d+2}{q_{1}}\bigr)(1-\theta)}N^{-\frac{\theta}{2}}\|f_{1}\|_{L^{2}(S^{d},\sigma)}\|\|_{L^{2}(S^{d},\sigma)}.$
We set $\alpha=\bigl(\frac{d+2}{q_{1}}-d\bigr)(1-\theta)+\frac{\theta}{2}$. Therefore we have proved Proposition 5.1. ∎
Remark 5.2.
As coined in [9], this is known as “distant caps interact weakly”. When $d=2$, $\frac{d+2}{d}=2$, there is no need to choose $q_{1}<2$. In this case, the proof above use Bourgain’s bilinear restriction estimate in [4, Lemma 111] and [16, Theorem 4.18]. We also remark that it provides an alternative way to the estimate in [9, Lemma 7.5], where Christ and the first author use the structure of convolution of the sphere surface measures.
Proposition 5.3 (Orthogonality II).
Let
$\limsup_{\nu\to\infty}\frac{r_{\nu}^{j}}{r_{\nu}^{k}}+\frac{r_{\nu}^{k}}{r_{\nu}^{j}}+\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}=\infty.$
We have that, for any $N\geq 1$, as $\nu\to\infty$,
(54) $\lim_{\nu\to\infty}\left\|\Bigl((r_{\nu}^{j})^{d/2}e^{\frac{it(r_{\nu}^{j})^{2}\Delta}{2}}\bigl(h_{\nu}((r_{\nu}^{j})^{2}t,\cdot)G_{\nu}^{j}\bigr)(r_{\nu}^{j}x)\Bigr)\Bigl((r_{\nu}^{k})^{d/2}e^{\frac{it(r_{\nu}^{k})^{2}\Delta}{2}}\bigl(h_{\nu}((r_{\nu}^{k})^{2}t,\cdot)G_{\nu}^{k}\bigr)(r_{\nu}^{k}x)\Bigr)\right\|_{L^{1+2/d}}=0$
for $G_{\nu}^{j}$, $G_{\nu}^{k}$ defined in (29).
Proof.
We assume that, up to subsequences, $\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\to\infty$ as $\nu\to\infty$. It is further split into two cases.
  • Case 1, $(r_{\nu}^{j},r_{\nu}^{k})\to(0,0)$. We change variables as follows, $r_{\nu}^{k}x\to x$ and $(r_{\nu}^{k})^{2}t\to t$. Then the norm on the left hand side of (54) equals
    (55) $\begin{split}&\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{d/2}\left\|e^{it\bigl(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\bigr)^{2}\Delta/2}\bigl(h_{\nu}(\bigl(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\bigr)^{2}t,y)G_{\nu}^{j}\bigr)(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}x)\Bigr)\,e^{\frac{it\Delta}{2}}\bigl(h_{\nu}(t,y)G_{\nu}^{k}\bigr)(x)\right\|_{L^{1+2/d}}\\ \quad&=\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{d/2}\left\|\int e^{i(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}x-x_{\nu}^{j})y+i\bigl((\frac{r_{\nu}^{j}}{r_{\nu}^{k}})^{2}t-t_{\nu}^{j}\bigr)\frac{\sqrt{1-|r_{\nu}^{j}y|^{2}}-1}{(r_{\nu}^{j})^{2}}}\dfrac{\phi^{j}(y)}{(1-|r_{\nu}^{j}y|^{2})^{1/4}}dy\right.\\ &\quad\quad\times\int\left.e^{i(x-x_{\nu}^{k})y+i\bigl(t-t_{\nu}^{k}\bigr)\frac{\sqrt{1-|r_{\nu}^{k}y|^{2}}-1}{(r_{\nu}^{k})^{2}}}\dfrac{\phi^{k}(y)}{(1-|r_{\nu}^{k}y|^{2})^{1/4}}dy\right\|_{L^{1+2/d}}.\end{split}$
    If $(r_{\nu}^{j},r_{\nu}^{k})\to(0,0)$, then $\frac{\sqrt{1-|r_{\nu}^{j}y|^{2}}-1}{(r_{\nu}^{j})^{2}}\to-\frac{|y|^{2}}{2}$ as $\nu\to\infty$; furthermore, we know that $\phi^{j}$ and $\phi^{k}$ can be taken to be smooth functions with compact supports in a ball of radius $1$ and uniform $L^{\infty}$ bound. These can be used via the Strichartz estimate for the quadratic surfaces [32] to show
    (56) $\left\|\int e^{ixy+it\frac{\sqrt{1-|r_{\nu}^{j}y|^{2}}-1}{(r_{\nu}^{j})^{2}}}\dfrac{\phi^{j}(y)}{(1-|y|^{2})^{1/4}}dy\right\|_{L^{2+4/d}(|x|+|t|>R)}\to 0,$
    as $R\to\infty$, uniformly in $\nu$. Indeed, by the change of variables, we have
    $(r_{\nu}^{j})^{-d}\|\int e^{i\frac{x}{r_{\nu}^{j}}y+i\frac{t}{(r_{\nu}^{j})^{2}}\sqrt{1-|y|^{2}}}\dfrac{\phi^{j}(\frac{y}{r_{\nu}^{j}})}{(1-|\frac{y}{r_{\nu}^{j}}|^{2})^{1/4}}dy\|_{L^{2+\frac{d}{4}}_{t,x}\big(|x|+|t|\geq R\bigr)}.$
    After changing variables $(r_{\nu}^{j})^{-1}x\to x$ and $(r_{\nu}^{j})^{-2}t\to t$, we see that
    $\begin{split}&\|\int e^{ix\cdot y+it\sqrt{1-|y|^{2}}}\dfrac{r_{\nu}^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})}{(1-|\frac{y}{r_{\nu}^{j}}|^{2})^{1/4}}\sqrt{1-|y|^{2}}\frac{dy}{\sqrt{1-|y||^{2}}}\|_{L^{2+\frac{d}{4}}_{t,x}\big(r_{\nu}^{j}|x|+(r_{\nu}^{j})^{2}|t|\geq R\bigr)}\\ &=\|\int e^{ix\cdot y+it\sqrt{1-|y|^{2}}}\dfrac{r_{\nu}^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})}{(1-|\frac{y}{r_{\nu}^{j}}|^{2})^{1/4}}\sqrt{1-|y|^{2}}d\sigma(y)\|_{L^{2+\frac{d}{4}}_{t,x}\big(r_{\nu}^{j}|x|+(r_{\nu}^{j})^{2}|t|\geq R\bigr)}.\end{split}$
    The $L^{2+\frac{4}{d}}_{t,x}$ norm is bounded by $\|\dfrac{r_{\nu}^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})}{(1-|\frac{y}{r_{\nu}^{j}}|^{2})^{1/4}}\sqrt{1-|y|^{2}}\|_{L^{2}(S^{d},\sigma)}$ by using the Strichartz estimate. If we expand this norm with the use of expressing the sphere surface measure $d\sigma$ in terms of the $y$-coordinates, and change variables, we see that is bounded by $\|\phi^{j}\|_{L^{2}}$. We observe that
    $\{r_{\nu}^{j}|x|+(r_{\nu}^{j})^{2}|t|\geq R\}\subset\{|x|+|t|\geq R/r_{\nu}^{j}\}\subset\{|x|+|t|\geq R\}.$
    since $r_{\nu}^{j}\in(0,\frac{1}{2})$. By the Strichartz estimate and the dominated convergence theorem, we see that it converges to zero as $\nu$ goes to infinity. Likewise for the $k$-th term. Hence we are led to consider the integration over the region $\Omega_{\nu}^{j}\cap\Omega_{\nu}^{k}$, where
    (57) $\begin{split}\Omega_{\nu}^{j}&:=\{\left|\frac{r_{\nu}^{j}}{r_{\nu}^{k}}x-x_{\nu}^{j}\right|+\left|(\frac{r_{\nu}^{j}}{r_{\nu}^{k}})^{2}t-t_{\nu}^{j}\right|\leq R\},\\ \Omega_{\nu}^{k}&:=\{\left|x-x_{\nu}^{k}\right|+\left|t-t_{\nu}^{k}\right|\leq R\}\end{split}$
    for a large $R>0$. Then continuing (55), we see that it is bounded by
    (58) $C_{R}\|\phi^{j}\|_{1}\|\phi^{k}\|_{1}\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{d/2}\min\{\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{-d},\,1\}\leq C_{R}\|\phi^{j}\|_{1}\|\phi^{k}\|_{1}\min\{\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{d/2},\,\left(\frac{r_{\nu}^{j}}{r_{\nu}^{k}}\right)^{-d/2}\}\to 0,$
    as $\nu\to\infty$.
  • Case 2, $(r_{\nu}^{j},r_{\nu}^{k})\to(c,0)$ for some $c>0$. It is similar to Case 1.
Then we may assume that $r_{\nu}^{j}\equiv r_{\nu}^{k}$ for all $\nu$. In this case, up to subsequences, we consider $\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}\to\infty$ as $\nu\to\infty$. This condition easily implies $r_{\nu}^{j}\to 0$ as $\nu\to\infty$.
As in Case 1 above, we estimate
(59) $\begin{split}&\left\|\int e^{i(x-x_{\nu}^{j})y+i((t-t_{\nu}^{j})\frac{\sqrt{1-|r_{\nu}^{j}y|^{2}}-1}{(r_{\nu}^{j})^{2}}}\phi^{j}(y)dy\,\int e^{i(x-x_{\nu}^{k})y+i\bigl(t-t_{\nu}^{k}\bigr)\frac{\sqrt{1-|r_{\nu}^{k}y|^{2}}-1}{(r_{\nu}^{k})^{2}}}\phi^{k}(y)dy\right\|_{L^{1+2/d}}\\ &\lesssim(r_{\nu}^{j})^{-d}\left\|\int e^{i\frac{x-x_{\nu}^{j}}{r_{\nu}^{j}}y+i\frac{t-t_{\nu}^{j}}{(r_{\nu}^{j})^{2}}\sqrt{1-|y|^{2}}}(r_{\nu}^{j})^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})dy\right.\\ &\quad\left.\times\int e^{i\frac{x-x_{\nu}^{k}}{r_{\nu}^{j}}y+i\frac{t-t_{\nu}^{k}}{(r_{\nu}^{j})^{2}}\sqrt{1-|y|^{2}}}(r_{\nu}^{j})^{-\frac{d}{2}}\phi^{k}(\frac{y}{r_{\nu}^{j}})\right\|_{L^{1+2/d}}\\ &\lesssim\left\|\int e^{i(x-(r_{\nu}^{j})^{-1}x_{\nu}^{j})\cdot y+i(t-(r_{\nu}^{j})^{-2}t_{\nu}^{j})\sqrt{1-|y|^{2}}}(r_{\nu}^{j})^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})dy\right.\\ &\quad\times\left.\int e^{i(x-(r_{\nu}^{j})^{-1}x_{\nu}^{k})\cdot y+i(t-(r_{\nu}^{j})^{-2}t_{\nu}^{k})\sqrt{1-|y|^{2}}}(r_{\nu}^{j})^{-\frac{d}{2}}\phi^{j}(\frac{y}{r_{\nu}^{j}})dy\right\|_{L^{1+2/d}}.\end{split}$
We recognize it is an adjoint bilinear restriction estimate for sphere, where the the lift-function-to-sphere $\phi^{j}(\frac{\cdot}{r_{\nu}^{j}})$ and $\phi^{k}(\frac{\cdot}{r_{\nu}^{j}})$ are supported in cap $\mathcal{C}_{\nu}^{j}$ and $\mathcal{C}_{\nu}^{k}$, respectively. Moreover we observe that
(60) $R_{\nu}:=\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}\to\infty,\text{ as }\nu\to\infty.$
Then by Lemma 5.1,
(61) $\begin{split}&\left\|\int e^{i(x-x_{\nu}^{j})y+i((t-t_{\nu}^{j})\frac{\sqrt{1-|r_{\nu}^{j}y|^{2}}-1}{(r_{\nu}^{j})^{2}}}\phi^{j}(y)dy\,\int e^{i(x-x_{\nu}^{k})y+i\bigl(t-t_{\nu}^{k}\bigr)\frac{\sqrt{1-|r_{\nu}^{k}y|^{2}}-1}{(r_{\nu}^{k})^{2}}}\phi^{k}(y)dy\right\|_{L^{1+\frac{2}{d}}}\\ &\leq CR_{\nu}^{-\alpha}\|\phi^{j}\|_{2}\|\phi^{k}\|_{2}\to 0.\end{split}$
Then by interpolation with the $L^{\infty}_{t,x}$ estimate, we see that the quantity in (59) converges to zero as $\nu\to\infty$. ∎
6. A synthesis
In this section, we will give a complete profile decomposition for $\{f_{\nu}\}_{\nu}\in L^{2}(S^{d})$, $d\geq 2$ where $\|f_{\nu}\|_{L^{2}(\Gamma,\sigma)}=1$. We will use the idea of a system of equations to express the decomposition.
  • 1.
    As in Proposition 3.5,
    (62) $f_{\nu}=\sum_{j=1}^{N}f_{\nu}^{j}+e^{N},$
    where each $f_{\nu}^{j}$ is supported on a cap $\mathcal{C}(z_{\nu}^{j},r_{\nu}^{j})$ with the other properties listed there.
  • 2.
    Motivated by the decomposition (22), as in Proposition 4.2,
    (63) $f_{\nu}^{j}(x)=\left((1-|\cdot|^{2})^{1/4}(r_{\nu}^{j})^{-\frac{d}{2}}g_{\nu}^{j}(\frac{\cdot}{r_{\nu}^{j}})\right)\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}(x),$
    where $g_{\nu}^{j}(y):=\dfrac{(r_{\nu}^{j})^{d/2}f_{\nu}^{j}\circ\Pi^{-1}_{H_{z_{\nu}^{j}}}L^{-1}_{z_{\nu}^{j}}(r_{\nu}^{j}y)}{(1-|r_{\nu}^{j}y|^{2})^{1/4}}$ and $L_{z},\Pi_{H_{z}}$ are defined in Definition 3.1.
  • 3.
    As in Proposition 4.2,
    (64) $g_{\nu}^{j}(y)=\sum_{\alpha=1}^{A_{j}}e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}(y)+e_{\nu}^{A_{j}}.$
  • 4.
    Combining all these, we have
    (65) $\begin{split}f_{\nu}(x)&=\sum_{j=1}^{N}\sum_{\alpha=1}^{A_{j}}\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}(x)\\ &\qquad\qquad\qquad+w_{\nu}^{N,A_{1},\cdots A_{N}}(x),\end{split}$
    where
    (66) $w_{\nu}^{N,A_{1},\cdots A_{N}}(x):=e_{\nu}^{N}+\sum_{j=1}^{N}\tilde{e}_{\nu}^{j,A_{j}}(x).$
    Here $\tilde{e}_{\nu}^{j,A_{j}}(x)=\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}e^{A_{j}}_{\nu}(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}(x).$
  • 5.
    The family $\Gamma_{\nu}^{j,\alpha}:=(r_{\nu}^{j},z_{\nu}^{j},x_{\nu}^{j,\alpha},t_{\nu}^{j,\alpha})_{1\leq j\leq N,\,1\leq\alpha\leq A_{j}}$ and $\Gamma_{\nu}^{k,\beta}:=(r_{\nu}^{k},z_{\nu}^{k},x_{\nu}^{k,\beta},t_{\nu}^{k,\beta})_{1\leq k\leq N,\,1\leq\beta\leq A_{k}}$ is pairwise orthogonal:
    (67) $\begin{split}&\forall j\neq k,\,\frac{r_{\nu}^{j}}{r_{\nu}^{k}}+\frac{r_{\nu}^{k}}{r_{\nu}^{j}}+\frac{|z_{\nu}^{j}-z_{\nu}^{k}|}{r_{\nu}^{j}}\to\infty,\text{ as }\nu\to\infty,\\ &\forall\alpha\neq\beta,|x_{\nu}^{j,\alpha}-x_{\nu}^{j,\beta}|+|t_{\nu}^{j,\alpha}-t_{\nu}^{j,\beta}|\to\infty,\text{ as }\nu\to\infty.\end{split}$
  • 6.
    Regarding the error terms, we have the following two estimates,
    (68) $\|w_{\nu}^{N,A_{1},\cdots,A_{N}}\|^{2}_{L^{2}(S^{d},\sigma)}:=\sum_{j=1}^{N}\|\tilde{e}_{\nu}^{j,A_{j}}\|^{2}_{L^{2}(S^{d},\sigma)}+\|e_{\nu}^{N}\|^{2}_{L^{2}(S^{d},\sigma)}=\sum_{j=1}^{N}\|e_{\nu}^{j,A_{j}}\|^{2}_{L^{2}(S^{d},\sigma)}+\|e_{\nu}^{N}\|^{2}_{L^{2}(S^{d},\sigma)},$
    (69) $\limsup_{\nu\to\infty}\|\widehat{w_{\nu}^{N,A_{1},\cdots,A_{N}}\sigma}\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\to 0,\text{ as }\min_{1\leq j\leq N}\{N,j+A_{j}\}\to\infty.$
    Equation (68) is due to the disjoint supports of the functions on the sphere. For the second, indeed, let $\delta>0$ be an arbitrary small number. Take $N_{0}$ such that for every $N\geq N_{0}$,
    (70) $\limsup_{\nu\to\infty}\|\widehat{e_{\nu}^{N}\sigma}\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\leq\frac{\delta}{3}.$
    For every $N\geq N_{0}$, there exists $B_{N}$ such that for $A_{j}\geq B_{N}$,
    (71) $\limsup_{\nu\to\infty}\|\widehat{\tilde{e}_{\nu}^{j,A_{j}}\sigma}\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\leq\frac{\delta}{3N}.$
    Then the reminder $w_{\nu}^{N,A_{1},\cdots,A_{N}}$ can be rewritten in the form
    (72) $w_{\nu}^{N,A_{1},\cdots,A_{N}}=e_{\nu}^{N}+\sum_{1\leq j\leq N}w_{\nu}^{j,A_{j}\vee B_{N}}+S_{\nu}^{N,A_{1},\cdots,A_{N}}.$
    Here $A_{j}\vee B_{N}=\max\{A_{j},B_{N}\}$ and
    (73) $S_{\nu}^{N,A_{1},\cdots,A_{N}}=\sum_{1\leq j\leq N,\atop A_{j}<B_{N}}(w_{\nu}^{j,A_{j}}-w_{\nu}^{j,B_{N}}),$
    which is a sum of profiles in form of (65). Therefore
    (74) $\|\widehat{w_{\nu}^{N}\sigma}\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\leq\frac{2\delta}{3}+\|\widehat{S_{\nu}^{N,A_{1},\cdots,A_{N}}}\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\leq\delta,$
    for sufficiently large $\nu$ due to Proposition 4.3 and Proposition 5.3.
To summarize the above, we have proved a complete profile decomposition.
Proposition 6.1.
Let $\{f_{\nu}\}_{1\leq\nu\leq\infty}\in L^{2}(S^{d})$ be a sequence of functions with $\|f_{\nu}\|_{L^{2}(S^{d},\sigma)}=1$. Then there exists a sequence of parameters
$\Gamma_{\nu}^{j,\alpha}:=(r_{\nu}^{j},z_{\nu}^{j},x_{\nu}^{j,\alpha},t_{\nu}^{j,\alpha})_{1\leq j<\infty,\,1\leq\alpha\leq A_{j}}\in(0,\infty)\times S^{d}\times\mathbb{R}^{d}\times\mathbb{R}$
satisfying (67) and $\{\phi^{j,\alpha}\}_{1\leq j<\infty,\,1\leq\alpha\leq A_{j}}\in L^{2}(\mathbb{R}^{d})$ such that
(75) $\begin{split}f_{\nu}(x)&=\sum_{j=1}^{N}\sum_{\alpha=1}^{A_{j}}\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}(x)\\ &\qquad\qquad\qquad+w_{\nu}^{N,A_{1},\cdots A_{N}}(x).\end{split}$
We have the following orthogonality identities.
(76) $\|f_{\nu}\|^{2}_{L^{2}(S^{d},\sigma)}=\sum_{j=1}^{N}\sum_{\alpha=1}^{A_{j}}\|\phi^{j,\alpha}\|^{2}_{L^{2}(\mathbb{R}^{d})}+\|w_{\nu}^{N,A_{1},\cdots,A_{N}}\|^{2}_{L^{2}(S^{d},\sigma)}+\|e_{\nu}^{N}\|^{2}_{L^{2}(S^{d},\sigma)},\text{ as }\nu\to\infty.$
(77) $\text{For }j\neq k,\left\|\mathcal{F}\left(\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}\sigma\right)\right$.
$\times\left.\mathcal{F}\left(\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{k})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{k,\beta}|y|^{2}}{2}}e^{-ix_{\nu}^{k,\beta}\cdot y}\phi^{k,\beta}\right)(\frac{\cdot}{r_{\nu}^{k}})\right]\circ L_{z_{\nu}^{k}}\circ\Pi_{H_{z_{\nu}^{k}}}\sigma\right)\right\|_{L^{1+\frac{2}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}\to 0,\text{ as }\nu\to\infty.$
(78) $\limsup_{\nu\to\infty}\|\mathcal{F}\left(w_{\nu}^{N,A_{1},\cdots,A_{N}}\sigma\right)\|_{L^{2+\frac{4}{d}}(\mathbb{R}\times\mathbb{R}^{d})}\to 0,\text{ as }\min_{1\leq j\leq N}\{N,j+A_{j}\}\to\infty.$
Here $\mathcal{F}(f\sigma)(x):=\widehat{f\sigma}(x)=\int_{S^{d}}e^{ix\cdot\xi}f(\xi)d\sigma(\xi)$.
One consequence of (77) is the following orthogonality result. The proof follows similar lines as in [2, Lemma 5.5].
Proposition 6.2.
Let $N$ and $\{A_{j}\}_{1\leq j\leq N}$ be given.
(79) $\left\|\sum_{j=1}^{N}\sum_{1\leq\alpha\leq A_{j}}\mathcal{F}\left(\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}\sigma\right)\right\|^{2+\frac{4}{d}}_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}$
$\leq\sum_{j=1}^{N}\sum_{1\leq\alpha\leq A_{j}}\left\|\mathcal{F}\left(\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}\sigma\right)\right\|^{2+\frac{4}{d}}_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}$
as $\nu\to\infty.$
7. Conditional existence
We assume that $\mathcal{R}>\mathcal{R}_{\textbf{P}}$. This is actually proved in [14] by make a real hypothesis that Gaussians are maximizers for the Strichartz inequality for the Schrödinger equation. Let us recall one notation from [14]. Define the optimal constant for the Strichartz inequality for the Schrödinger equation, we see that
(80) $\mathcal{R}_{\mathbf{P}}=(2\pi)^{-\frac{d+2}{d}}\dfrac{\int_{\mathbb{R}^{d+1}}|e^{it\Delta/2}\phi_{G}(x)|^{2+\frac{4}{d}}dxdt}{\|\phi_{G}\|^{2+\frac{d}{4}}_{L^{2}(\mathbb{R})}},$
where $\phi_{G}(x)=e^{-|x|^{2}/2}$. The [14, Lemma 7.2] shows that
$\mathcal{R}>\mathcal{R}_{\mathbf{P}}.$
We first note that when $r_{\nu}\to 0$, for a compactly supported and smooth function $\phi^{j}$,
(81) $\int e^{ixy-\frac{it|y|^{2}}{2}}e^{it\bigl(\frac{\sqrt{1-r_{\nu}^{2}|y|^{2}}-1}{r_{\nu}^{2}}+\frac{|y|^{2}}{2}\bigr)}\phi^{j}(y)dy\to e^{it\Delta/2}\phi^{j},$
in the $L^{2+\frac{d}{4}}_{t,x}$ Strichartz norm sense, and in the pointwise sense. This follows from the standard stationary phase estimates as in (46), (47) and (48) and the dominated convergence theorem.
Now we begin our argument for the conditional existence. By Propositions 6.1 and 6.2,
(82) $\limsup_{\nu\to\infty}\|\mathcal{F}(f_{\nu}\sigma)\|_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}^{2+4/d}$
$\leq\sum_{j=1}^{N}\sum_{1\leq\alpha\leq A_{j}}\limsup_{\nu\to\infty}\left\|\mathcal{F}\left(\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu}^{j})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}^{j,\alpha}|y|^{2}}{2}}e^{-ix_{\nu}^{j,\alpha}\cdot y}\phi^{j,\alpha}\right)(\frac{\cdot}{r_{\nu}^{j}})\right]\circ L_{z_{\nu}^{j}}\circ\Pi_{H_{z_{\nu}^{j}}}\sigma\right)\right\|^{2+\frac{4}{d}}_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}$
$\leq\sum_{j}\limsup_{\nu\to\infty}\left\|\int e^{i(x-x_{\nu}^{j})y-\frac{i(t-t_{\nu}^{j})|y|^{2}}{2}}e^{i(t-t_{\nu}^{j})\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\phi^{j}(y)dy\right\|_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}^{2+4/d}$
$=\sum_{j}\limsup_{\nu\to\infty}\left\|\int e^{ixy-\frac{it|y|^{2}}{2}}e^{it\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\phi^{j}(y)dy\right\|_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}^{2+4/d}.$
Here we are abusing the notations of summations when passing from the second line to the third line and also increase the sum to the infinite terms. We consider two cases. If $r_{\nu}^{j}\to 0$ for some $j$,
(83) $\begin{split}&\limsup_{\nu\to\infty}\left\|\int e^{ixy-\frac{it|y|^{2}}{2}}e^{it\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\phi^{j}(y)dy\right\|_{L^{2+\frac{4}{d}}_{t,x}(\mathbb{R}\times\mathbb{R}^{d})}^{2+4/d}\\ &=\|e^{\frac{it\Delta}{2}}\phi^{j}\|_{2+4/d}^{2+4/d}\leq\mathcal{R}^{2+4/d}_{\textbf{P}}\|\phi^{j}\|_{2}^{2+4/d}\\ &\leq\mathcal{R}^{2+4/d}_{\textbf{P}}\|\phi^{j}\|_{2}^{2+4/d}<\mathcal{R}^{2+\frac{4}{d}}\|\phi^{j}\|^{2+\frac{4}{d}}_{L^{2}},\end{split}$
by the assumption that $\mathcal{R}>\mathcal{R}_{\mathbf{P}}$. On the other hand, if $r_{\nu}^{j}\to r^{j}$ for some $j$, by the Strichartz estimate,
(84) $\begin{split}&\limsup_{\nu\to\infty}\left\|\int e^{ixy-\frac{it|y|^{2}}{2}}e^{it\bigl(\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}-1}{(r_{\nu}^{j})^{2}}+\frac{|y|^{2}}{2}\bigr)}\phi^{j}(y)dy\right\|_{2+4/d}^{2+4/d}\\ &=\limsup_{\nu\to\infty}\left\|\int e^{ixy+it\frac{\sqrt{1-(r_{\nu}^{j})^{2}|y|^{2}}}{(r_{\nu}^{j})^{2}}}\phi^{j}(y)dy\right\|_{2+4/d}^{2+4/d}\\ &=\limsup_{\nu\to\infty}\left\|\int e^{ixy+it\sqrt{1-|y|^{2}}}\left((r_{\nu}^{j})^{-d/2}\phi^{j}((r_{\nu}^{j})^{-1}y)\right)dy\right\|_{2+4/d}^{2+4/d}\\ &=\limsup_{\nu\to\infty}\left\|\int e^{ixy+it\sqrt{1-|y|^{2}}}\left((r_{\nu}^{j})^{-d/2}\phi^{j}((r_{\nu}^{j})^{-1}y)\sqrt{1-|y|^{2}}\right)\frac{dy}{\sqrt{1-|y|^{2}}}\right\|_{2+4/d}^{2+4/d}\\ &=\limsup_{\nu\to\infty}\left\|\int e^{ixy+it\sqrt{1-|y|^{2}}}\left((r^{j})^{-d/2}\phi^{j}((r^{j})^{-1}y)\sqrt{1-|y|^{2}}\right)\frac{dy}{\sqrt{1-|y|^{2}}}\right\|_{2+4/d}^{2+4/d}\\ &=\mathcal{R}^{2+4/d}\|(r^{j})^{-d/2}\phi^{j}((r^{j})^{-1}y)\sqrt{1-|y|^{2}}\|_{L^{2}(\sigma)}^{2+4/d}\\ &=\mathcal{R}^{2+4/d}\|\phi^{j}(y)(1-|r^{j}y|^{2})^{1/4}\|_{L^{2}}^{2+4/d}\leq\mathcal{R}^{2+4/d}\|\phi^{j}\|_{L^{2}}^{2+4/d}.\end{split}$
Here we have used the Tomas-Stein inequality for the sphere and the dominated convergence theorem in passing the fourth line to the fifth line above. So combining (83) and (84), we see that
[…]
Then $\mathcal{R}^{2+4/d}=\mathcal{R}^{2+4/d}$ forces all the inequalities above to be equal. On the other hand, because $\sum_{j}\|\phi^{j}\|_{L^{2}}^{2}\leq 1$, there will be only one $j$ left from the sharpness of embedding of $\ell^{1+2/d}$ into $\ell^{1}$. If $r_{\nu}^{j}\to r^{j}>0$ for this $j$, we return to (82) and (84), and obtain that $(r^{j})^{-d/2}\phi((r^{j})^{-1}y)\sqrt{1-|y|^{2}}$ is an extremal as desired, which is smooth. If $r_{\nu}^{j}\to 0$ for this $j$, then we return to (82) and (83), and obtain a contradiction to the assumption $\mathcal{R}>\mathcal{R}_{\textbf{P}}$. Indeed, because there is one term $\phi^{j}$ left,
(86) $f_{\nu}=\left[(1-|\cdot|^{2})^{\frac{1}{4}}(r_{\nu})^{-\frac{d}{2}}\left(e^{\frac{it_{\nu}|y|^{2}}{2}}e^{-ix_{\nu}\cdot y}\phi^{j}\right)(\frac{\cdot}{r_{\nu}})\right]\circ L_{z_{\nu}}\circ\Pi_{H_{z_{\nu}}},$
(87) $\limsup_{\nu\to\infty}\|\mathcal{F}(f_{\nu}\sigma)\|^{2+\frac{4}{d}}_{L^{2+\frac{4}{d}}_{t,x}}=\mathcal{R}^{2+\frac{4}{d}}\limsup_{\nu\to\infty}\|f_{\nu}\|^{2+\frac{4}{d}}_{L^{2}(S^{d},\sigma)}=\mathcal{R}^{2+\frac{4}{d}}\|\phi^{j}\|^{2+\frac{4}{d}}_{L^{2}(\mathbb{R}^{d})}.$